### v2

# Complex Domain Approach for Reversible Data Hiding and Homomorphic Encryption: General Framework and Application to Dispersed Data

Abstract
Ensuring the trustworthiness of data from distributed and resource-constrained environments, such as Wireless Sensor Networks or IoT devices, is critical. Existing Reversible Data Hiding (RDH) methods for scalar data suffer from low embedding capacity and poor intrinsic mixing between host data and watermark. This paper introduces Hiding in the Imaginary Domain with Data Encryption (H[i][i]dden), a novel framework based on complex number arithmetic for simultaneous information embedding and encryption. The H[i][i]dden framework offers perfect reversibility, in-principle unlimited watermark size, and intrinsic data-watermark mixing. The paper further introduces two protocols: H[i][i]dden-EG, for joint reversible data hiding and encryption, and H[i][i]dden-AggP, for privacy-preserving aggregation of watermarked data, based on partially homomorphic encryption. These protocols provide efficient and resilient solutions for data integrity, provenance and confidentiality, serving as a foundation for new schemes based on the algebraic properties of the complex domain.
Keywords: Reversible Data Hiding, Complex Domain, Gaussian Integers, Homomorphic Encryption, Data Aggregation, Dispersed Data, Paillier, ElGamal, Data Integrity
1 Introduction
The ubiquitous generation and pervasive use of digital data stand as defining characteristics of the modern era. From vast networks of Internet of Things (IoT) devices and industrial sensors to intricate smart city infrastructures and critical healthcare systems, data fuels decision-making, automates processes, and enables unprecedented levels of connectivity. This reliance on data, often collected from distributed and potentially vulnerable environments, is critically linked to its trustworthiness (Yan et al., 2014). Ensuring the confidentiality, integrity, authenticity, and privacy of information has become a priority, as compromised data can lead to erroneous conclusions, system failures, and severe breaches of privacy or security (Liang et al., 2016).
Within this complex digital landscape, the concept of Reversible Data Hiding (RDH) (Shi et al., 2016; Kumar et al., 2022) emerges as a specialized and increasingly popular approach to data security. RDH involves embedding secret information (a “watermark”) directly into a host data item in such a way that, after extraction of the watermark, the original host data can be perfectly and exactly restored to its original state. This unique property, combining data embedding with guaranteed reversibility, makes RDH appropriate in sensitive applications such as medical imaging, legal documentation, and military intelligence, where any permanent alteration of the original data is unacceptable. Unlike traditional data encryption, which primarily ensures confidentiality, RDH focuses on providing intrinsic integrity, authenticity, and provenance, within the data itself.
In the context of individual, scalar, or time-series numerical data, such as those originating from Wireless Sensor Networks (WSNs) (Shi and Xiao, 2013), IoT devices (Rahman et al., 2021), or critical financial and medical monitoring systems (Bhalerao et al., 2019), RDH offers unique advantages. Foremost among these is its ability to ensure robust data integrity and authenticity at the granular level of single readings or sequential observations. By embedding a watermark that can be extracted and verified, receivers (e.g., a Data Collector in a WSN) can confirm that the data originated from a legitimate source and has not been tampered with in transit. This inherent capability to facilitate data provenance and traceability is particularly valuable, allowing for forensic analysis of data origin and processing paths in distributed and often resource-constrained environments.
Beyond these general benefits, RDH offers distinct advantages compared to other commonly employed security primitives in certain contexts. Unlike cryptographic hashing or digital signatures, which typically append a separate authentication tag or signature to the data, RDH embeds the integrity verification mechanism directly within the data itself. This can simplify data handling and ensure the integrity proof travels jointly with the content. While conventional encryption ensures confidentiality, it does not directly provide data integrity; and Physical Unclonable Functions (PUFs) (Gao et al., 2020) focus on device authentication rather than data integrity. RDH uniquely combines content-dependent integrity (through watermark verification) with the guarantee of perfect original data restoration, a feature not offered by one-way hash functions or irreversible authentication tags. This intrinsic embedding can be particularly beneficial in constrained environments or specialized data formats where appending additional metadata is undesirable or challenging.
From an implementation perspective, basic RDH techniques can often be quite lightweight, requiring minimal computational resources. This aspect is particularly useful for deployment on resource-constrained devices like low-power sensors. However, it is worth pointing out that this “lightweight” advantage primarily applies to simple, non-encrypted RDH schemes. When RDH is combined with advanced cryptographic primitives, such as the homomorphic encryption schemes explored in this paper, the computational overhead significantly increases, with cryptographic operations becoming the dominant factor. Nevertheless, the enhanced security and privacy offered by such combination often justify this increased complexity, addressing critical requirements for secure and trustworthy data ecosystems.
Despite these advantages, a significant limitation of existing RDH techniques for individual or dispersed data, particularly in high-security contexts, is their low embedding capacity. Typically, these methods allow for the embedding of only 1 bit per data sample or even less (as is often the case with difference expansion methods). This low capacity severely limits the granularity of tampering detection and the strength of integrity guarantees for independent, single data items. For instance, if only 1 bit is embedded in a single measurement, the probability of an attacker randomly guessing the correct bit is 1/2, offering very weak protection. While the security scales with the number of samples, making the probability of deceiving detection negligible for a sufficient number of measurements $M$ (e.g., $2^{-M}$), this aggregate protection is not robust in situations where the frequency of measurements or the total number of independent data points is small. This highlights a gap in current RDH capabilities for truly strong, individual data item integrity in such sparse or low-frequency scenarios.
In a parallel development within cryptography, Homomorphic Encryption (HE) (Acar et al., 2018) has emerged as a transformative technology for data privacy (Iezzi, 2020). HE schemes allow computations to be performed directly on encrypted data without ever needing to decrypt it (Zhang et al., 2020). The output of such a computation is an encrypted result that, when decrypted, is identical to the result of the same computation performed on the unencrypted data. This property enables the processing of sensitive information by untrusted third parties (such as cloud servers) while ensuring that the data’s confidentiality remains fully protected throughout its lifecycle.
While HE is broadly applied to solve privacy challenges in cloud computing, its principles can be integrated with other security primitives to achieve more complex security goals. For instance, combining data integrity mechanisms like RDH with HE’s ability to operate on encrypted data allows for the creation of protocols that simultaneously guarantee confidentiality, authenticity, and data integrity. This powerful synergy is especially relevant in systems that require both privacy during data processing and proof of data origin and integrity, setting the stage for more robust and trustworthy data ecosystems.
1.1 Plan and Contributions of the Paper
This paper introduces a novel framework for RDH and encryption, taking advantage of the algebraic properties of the complex domain (Gaussian integers). Termed Hiding in the Imaginary Domain with Data Encryption (H[i][i]dden), the proposed approach fundamentally addresses key limitations of existing RDH techniques when applied to individual, scalar, or dispersed data. Unlike methods that merely append or isolate hidden bits in low-significance positions (e.g., modulo-based schemes), H[i][i]dden uses complex number arithmetic to mix the original data and the watermark, providing a richer embedding space and assuring a larger tampering detection capacity.
The main contributions of the paper are summarized below:
  • A novel core complex-domain reversible information hiding scheme for individual data, offering intrinsic data-watermark mixing, perfect reversibility and, in principle, unlimited watermark size.
  • A distinct application of the core H[i][i]dden embedding scheme to aggregated data, which prepares individual data items for privacy-preserving and efficient summation of watermarked data while retaining extractability.
  • A detailed novel extension of ElGamal encryption to the Gaussian integers, serving as a foundational cryptographic primitive for the complex domain.
  • The H[i][i]dden-EG protocol, providing joint reversible information hiding and encryption specifically for individual data points, making use of the extended Gaussian ElGamal scheme.
  • A novel component-wise application of Paillier encryption to complex-valued data, establishing a privacy-preserving additive homomorphism for secure aggregation.
  • The H[i][i]dden-AggP protocol, enabling joint reversible information hiding, encryption, and privacy-preserving data aggregation from multiple sources, built upon the component-wise Paillier application.
Through these contributions, the paper demonstrates a new paradigm for reversibly embedding authenticating information directly into sensitive data, providing strong cryptographic protection while maintaining the relevant property of perfect host data reversibility.
The rest of the paper is organized as follows. Section 2 provides the necessary background on RDH and HE. Section 3 details the proposed complex-domain reversible watermarking scheme for both individual and aggregated data. Section 4 presents the extension of ElGamal and Paillier encryption to the complex domain. Section 5 details the H[i][i]dden-EG and H[i][i]dden-AggP protocols for joint watermarking and encryption/aggregation, respectively. A comprehensive formal security and privacy analysis of the proposed protocols is provided in Section 6. Finally, Section 7 concludes the paper and outlines future research directions.
2 Background
This section provides the foundational concepts and related work that are essential for understanding the proposed framework. The core of this work lies at the intersection of two distinct but complementary fields of research: RDH and HE. While RDH offers powerful mechanisms for embedding information within data for integrity and provenance, HE provides the unique capability to perform computations on encrypted data, thereby ensuring confidentiality. The following sections detail the relevant background on these two key areas.
2.1 Reversible Data Hiding
Reversible Data Hiding (Shi et al., 2016; Kumar et al., 2022) is a specialized field within information hiding that focuses on embedding secret information into host data in such a way that the original host data can be perfectly and exactly recovered after the hidden data (watermark) has been extracted. This property is essential in applications that do not allow the permanent alteration of the original content.
Historically, RDH techniques have been extensively developed and applied primarily to multimedia content, including digital images, audio, and video (Shi et al., 2016). Methods like Least Significant Bit (LSB) modification (when combined with original LSB storage) (Celik et al., 2002), Difference Expansion (DE) (Tian, 2003), Histogram Shifting (HS) (Ni et al., 2006), and Prediction-Error Expansion (PEE) (Ou et al., 2013) use the inherent statistical redundancy and the human perceptual system’s tolerance for slight alterations in these rich data formats. These approaches can embed information in an imperceptible mmanner while ensuring perfect reversibility, meaning the host data is faithfully restored to its original state after hidden data extraction.
In contrast, the application of RDH to individual, scalar, or time-series numerical data (e.g., sensor readings, financial figures, or simple numerical sequences) presents distinct and more challenging problems. Unlike multimedia, such data typically exhibits very low inherent redundancy, is highly sensitive to even minor numerical changes (as there is no “perceptual” domain to hide in), and has a very limited host capacity. This scarcity of “hiding space” and the demand for exact numerical preservation make traditional RDH techniques, often designed for media, largely unsuitable.
For this class of individual or dispersed data, existing reversible embedding methods tend to fall into a few categories. One common approach is Modulo or Remainder-based embedding, where the secret information $w$ (an integer in range $0,\ldots,q-1$) is embedded into a host integer $d$ using an operation like $d^{\prime}=q\cdot d+w$ (Kabir et al., 2025). A common special case involves embedding a single bit ($w\in\{0,1\}$) by setting $q=2$, resulting in $d^{\prime}=2d+w$ (Kabir et al., 2024). Upon extraction, $w$ is simply $d^{\prime}\,(\operatorname{\mathrm{mod}}q)$ and $d$ is $\lfloor d^{\prime}/q\rfloor$. When $q$ is a power of 2 ($q=2^{B}$), these extraction operations can be efficiently performed using bitwise logic (e.g., $w=d^{\prime}\text{ \& }(q-1)$ and $d=d^{\prime}\gg B$).
While perfectly reversible and simple, a significant characteristic of this method is that the host data’s magnitude is directly scaled or inflated by a factor of $q$, making the numerical modification immediately obvious and leading to a poor mixing between the original data and the watermark.
While modulo-based embedding places the watermark in a very distinct (e.g., least significant) position within the embedded data, one might consider more intricate bit-level operations to “mix” the bits of the original data and the watermark more thoroughly within the resulting embedded value. Such approaches would likely involve complex sequences of bitwise operations, permutations, or scrambling algorithms. However, these methods, even if designed to be perfectly reversible, tend to lack mathematical elegance and direct interpretability. The resulting embedded data would exhibit a less intuitive relationship to its original host and watermark, making such bit-mixing techniques generally less desirable compared to approaches that achieve blending through inherent algebraic properties.
Another relevant approach for sequential data is Difference Expansion (DE), or its more general form, Prediction-Error Expansion (Shi and Xiao, 2013). These methods exploit relationships between consecutive data samples by computing prediction errors (e.g., $e_{k}=d_{k}-d_{k-1}$) and then embedding information by expanding these errors.
These existing methods for individual/scalar data often suffer from relatively low embedding capacity. This highlights the need for novel approaches that can integrate data hiding more subtly or within a more complex framework, especially when combined with other security operations like encryption.
2.2 Homomorphic Encryption
Homomorphic Encryption (Acar et al., 2018) is a unique and powerful class of cryptographic schemes that allows a user to perform computations on encrypted data without first having to decrypt it. The result of these operations is an encrypted output which, when decrypted, is the same as if the operations had been performed on the original, unencrypted data. This remarkable capability is useful for maintaining data privacy and confidentiality when computations are outsourced to untrusted third parties, such as cloud services or remote data aggregators.
HE schemes are typically classified based on the types and number of operations they support. Partially Homomorphic Encryption (PHE) schemes, such as (Cominetti and Simplicio, 2020), support an unlimited number of a single type of operation, either addition or multiplication. Somewhat Homomorphic Encryption (SHE) schemes, which include the pioneering work by Boneh-Goh-Nissim (BGN) in (Boneh et al., 2005), can perform a limited number of both additions and multiplications. Finally, Fully Homomorphic Encryption (FHE) schemes, such as Ring Learning with Errors (RLWE) (Brakerski and Vaikuntanathan, 2011), support an unlimited number of both operations, allowing for arbitrary computations on encrypted data. In this paper, two classical PHE schemes, ElGamal (1985) and Paillier (1999), are applied for their distinct homomorphic properties that enable the proposed protocols for both individual data encryption and privacy-preserving data aggregation.
2.2.1 ElGamal Encryption
ElGamal encryption (ElGamal, 1985) is an asymmetric key encryption algorithm based on the Diffie-Hellman key exchange scheme. Its security relies on the difficulty of the discrete logarithm problem (DLP) in the multiplicative group of integers modulo a prime number.
The cryptosystem works as follows. Let $p$ be a prime number and $g$ be a generator of the cyclic multiplicative group $\mathbb{Z}_{p}^{*}$, often also denoted as $\left({\mathbb{Z}/p\mathbb{Z}}\right)^{\times}$. The ElGamal encryption scheme consists of three algorithms:
  • Key generation: The key generation process involves selecting a random integer $a$ such that $1\leq a\leq p-1$ as the private key. Then, $K=g^{a}\,(\operatorname{\mathrm{mod}}p)$ is computed. The public key is formed as the triple $\mathbf{K}=(p,g,K)$.
  • Encryption: To encrypt a message $m\in\mathbb{Z}_{p}^{*}$, a random integer $b$, such that $1\leq b\leq p-1$, is selected as the ephemeral key. The ciphertext is then computed as $y_{1}=g^{b}\,(\operatorname{\mathrm{mod}}p)$ and $y_{2}=m\cdot K^{b}\,(\operatorname{\mathrm{mod}}p)$, resulting in the pair $(y_{1},y_{2})$. The encryption function can be denoted as follows:
    $(y_{1},y_{2})=\mathrm{Enc}_{\mathbf{K}}(m,b)=\left(g^{b}\,(\operatorname{\mathrm{mod}}p),m\cdot K^{b}\,(\operatorname{\mathrm{mod}}p)\right).$
  • Decryption: The decryption of the ciphertext $(y_{1},y_{2})$ using the private key $a$ involves computing the shared secret $s=y_{1}^{a}\,(\operatorname{\mathrm{mod}}p)$ and then decrypting the message as $m=y_{2}\cdot s^{-1}\,(\operatorname{\mathrm{mod}}p)$. This decryption function can be formally denoted as follows:
    $\mathrm{Dec}_{a}(y_{1},y_{2})=y_{2}\cdot(y_{1}^{a})^{-1}\,(\operatorname{\mathrm{mod}}p).$
The security of ElGamal encryption relies on the difficulty of the DLP in $\mathbb{Z}_{p}^{*}$. If an attacker can efficiently solve the DLP, they can compute the private key $a$ from the public key $K$ and then decrypt any ciphertext. However, for sufficiently large prime numbers $p$, the DLP is believed to be computationally infeasible.
ElGamal encryption exhibits a homomorphic property under component-wise multiplication of ciphertexts. This means that given two ciphertexts $\mathrm{Enc}_{\mathbf{K}}(m_{1},b_{1})=(y_{1},y_{2})$ and $\mathrm{Enc}_{\mathbf{K}}(m_{2},b_{2})=(y^{\prime}_{1},y^{\prime}_{2})$, multiplying them component-wise as $(y_{1}\cdot y^{\prime}_{1},y_{2}\cdot y^{\prime}_{2})$ results in a ciphertext that decrypts to the product of the original plaintexts:
$\mathrm{Dec}_{a}\left(\mathrm{Enc}_{\mathbf{K}}(m_{1},b_{1})\cdot\mathrm{Enc}_{\mathbf{K}}(m_{2},b_{2})\right)=m_{1}\cdot m_{2}\,(\operatorname{\mathrm{mod}}p).$
This property has applications in various cryptographic protocols and it is exploited to combine reversible watermarking and encryption in this paper.
2.2.2 Paillier Encryption
Paillier (1999) is a probabilistic asymmetric encryption algorithm. Unlike ElGamal, which is homomorphic with respect to multiplication, Paillier encryption is additively homomorphic, meaning that sums of plaintexts can be computed from their ciphertexts.
The cryptosystem works as follows. It operates in the group $\mathbb{Z}_{n^{2}}^{*}$, where $n$ is a product of two large prime numbers. The Paillier encryption scheme consists of three algorithms:
  • Key generation:
    1. 1.
      Select two large random prime numbers $p$ and $q$ such that $p\neq q$ and $\mathrm{gcd}(pq,(p-1)(q-1))=1$.
    2. 2.
      Compute $n=pq$ and $L_{P}=\mathrm{lcm}(p-1,q-1)$, where $\mathrm{lcm}(\cdot)$” stands for the least common multiple.
    3. 3.
      Select a random integer $g\in\mathbb{Z}_{n^{2}}^{*}$ such that $g$ has order a multiple of $n$ modulo $n^{2}$. A common choice is $g=n+1$.
    4. 4.
      Compute $M_{P}=(L(g^{L_{P}}\,(\operatorname{\mathrm{mod}}n^{2})))^{-1}\,(\operatorname{\mathrm{mod}}n)$, where the function $L(x)$ is defined as
      $L(x)=\frac{x-1}{n}.$
    The public key is formed as the pair $\mathsf{K_{pub}}=(n,g)$, and the private key is $\mathsf{K_{priv}}=(L_{P},M_{P})$.
  • Encryption: To encrypt a message $m$, where $0\leq m<n$, a random integer $r$ is selected such that $0<r<n$ and $\mathrm{gcd}(r,n)=1$. The ciphertext $c$ is then computed as:
    $c=g^{m}\cdot r^{n}\,(\operatorname{\mathrm{mod}}n^{2}).$
    The encryption function can be denoted as
    $\mathrm{Enc}_{\mathsf{K_{pub}}}(m,r)=g^{m}\cdot r^{n}\,(\operatorname{\mathrm{mod}}n^{2}).$
  • Decryption: The decryption of a ciphertext $c$ using the private key $(L_{P},M_{P})$ involves computing:
    $m=L(c^{L_{P}}\,(\operatorname{\mathrm{mod}}n^{2}))\cdot M_{P}\,(\operatorname{\mathrm{mod}}n).$
    This decryption function can be formally denoted as $\mathrm{Dec}_{\mathsf{K_{priv}}}(c)=L(c^{L_{P}}\,(\operatorname{\mathrm{mod}}n^{2}))\cdot M_{P}\,(\operatorname{\mathrm{mod}}n)$.
The security of Paillier encryption relies on the computational difficulty of the composite residuosity problem. This problem is considered hard for sufficiently large $n$, making it computationally infeasible for an attacker to recover the plaintext $m$ from a ciphertext $c$ without knowing the private key $(L_{P},M_{P})$.
A core feature of Paillier encryption is its additive homomorphic property. This means that operations can be performed on ciphertexts such that the decryption of the result corresponds to the addition of the original plaintexts.
  1. 1.
    Addition of encrypted numbers: Given two ciphertexts $c_{1}=\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{1},r_{1})$ and $c_{2}=\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{2},r_{2})$, their product modulo $n^{2}$ decrypts to the sum of their plaintexts:
    $\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{1}\cdot c_{2}\,(\operatorname{\mathrm{mod}}n^{2}))=(m_{1}+m_{2})\,(\operatorname{\mathrm{mod}}n).$
  2. 2.
    Multiplication by a plaintext constant: Given a ciphertext $c_{1}=\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{1},r_{1})$ and a plaintext constant $k$, the ciphertext raised to the power of $k$ modulo $n^{2}$ decrypts to the product of the plaintext and the constant:
    $\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{1}^{k}\,(\operatorname{\mathrm{mod}}n^{2}))=k\cdot m_{1}\,(\operatorname{\mathrm{mod}}n).$
This additive homomorphism is particularly useful for applications requiring privacy-preserving aggregation, where sums of data points need to be computed without revealing individual values. In the context of this paper, it will be exploited for secure data aggregation in a new protocol.
3 Reversible Watermarking Approach on the Complex Domain
This section presents a reversible watermarking scheme that can be used for isolated data measurements (e.g., readings of a smart meter or sensor data).
3.1 Basic Definitions
First, let us define the different elements of the reversible watermarking scheme.
Definition 1 (Fine-grained original data).
$d_{j,k}$ is the scalar data produced by the $j$-th device (e.g., a sensor) in the $k$-th time sample (time $t_{k}$).
Without loss of generality, we assume that the original data is an integer: $d_{j,k}\in\mathbb{Z}$, since it is represented on a digital device and can always be mapped to an integer (even if it is a floating-point number).
Definition 2 (Watermark).
$w_{k}$ is the watermark to be embedded at time $k$ in the fine-grained data $d_{j,k}$. Similarly, we assume $w_{k}\in\mathbb{Z}$.
Remark 1.
The only assumption we are making with respect to the watermark $w_{k}$ is to be an integer. Hence, any limitations in a practical application (e.g., a given number of bits) would be caused by the implementation details, such as the number of bits used for representing integers in a particular programming language. $\triangle$
Definition 3 (Reversible embedding and blind extraction functions).
The embedding process is carried out using an embedding function $E_{w}$ in the following way:
$d^{\prime}_{j,k}=E_{w}(d_{j,k},w_{k},K_{w}),$
where $K_{w}$ is a watermarking key required for both the embedding and extraction processes. The reversible extraction or recovery function $R_{w}$ allows recovery of both the original data and the watermark, as follows:
$\left[d_{j,k},w_{k}\right]^{\mathrm{T}}=R_{w}(d^{\prime}_{j,k},K_{w}),$
for the watermarked data $d^{\prime}_{j,k}$, where the superscript $\mathrm{T}$” denotes the transposition operator.
Note that the extraction is assumed to be blind, since the original data is not used in the function.
Remark 2.
Unlike reversible data hiding for multimedia content or highly redundant data —where source redundancy allows for techniques such as difference expansion, histogram shifting, or prediction error expansion to hide information without necessarily increasing carrier size— embedding a watermark $w_{k}$ in a single scalar $d_{j,k}$ such that both can be perfectly recovered requires data expansion. For instance, as seen in Kabir et al. (2025), embedding a $B$-bit watermark using an operation like $d^{\prime}_{j,k}=2^{B}\cdot d_{j,k}+w_{k}$ increases the bit-length of the original data by exactly $B$ bits. From an information-theoretic perspective, if the host data consists of $A$ bits and the watermark comprises $B$ bits, and both are independent and occupy their full possible ranges, the combined state space consists of $2^{A+B}$ distinct values. In such cases, the result must be encoded with at least $A+B$ bits for perfect recovery. While “slack” in a data’s restricted range (e.g., values $d_{j,k}\in[101,200]$ encoded in a 1-byte container) can theoretically host bits without expansion, scalar reversible embedding of independent, full-range variables fundamentally requires increasing the bit-length of the transmitted values. $\triangle$
Definition 4 (Aggregated data and watermarked aggregated data).
The aggregated data $S_{k}$, for the devices $j=1,2,\dots,N$ in the $k$-th time sample is defined as follows:
$S_{k}=\sum_{j=1}^{N}d_{j,k}.$
Similarly, the aggregated watermarked data $S^{\prime}_{k}$ can be defined as detailed below:
$S^{\prime}_{k}=\sum_{j=1}^{N}d^{\prime}_{j,k}=\sum_{j=1}^{N}E_{w}(d_{j,k},w_{k},K_{w}).$
Note that the watermark embedded in the different devices is the same ($w_{k}$), although it can vary at every $t_{k}$.
Definition 5 (Reversible blind aggregated data and watermark extraction).
An aggregated data and watermark extraction function $R^{\mathrm{Agg}}_{w}$ allows extracting both $S_{k}$ and $w_{k}$ from $S^{\prime}_{k}$ as follows:
$\left[S_{k},w_{k}\right]^{\mathrm{T}}=R^{\mathrm{Agg}}_{w}(S^{\prime}_{k},K_{w},[N]),$
where “[N][N]” indicates that the extraction function may require, as an additional parameter, the number of data samples involved in the aggregation.
3.2 Proposed Reversible Watermarking Scheme
This section details the embedding, extraction, and aggregated extraction functions for the proposed watermarking scheme.
First, let the watemarking key $K_{w}$ be a Gaussian integer $\lambda_{k}=a_{k}+ib_{k}$, where $i=\sqrt{-1}$ is the positive imaginary unit, $a_{k},b_{k}\in\mathbb{Z}$, and $a_{k},b_{k}\neq 0$. The last condition guarantees that $\lambda_{k}$ is not either just real or imaginary, and also that $\lambda_{k}\neq 0$. Thus, there exists a complex number $\lambda_{k}^{-1}$ such that $\lambda_{k}\lambda_{k}^{-1}=1$, with:
$\lambda_{k}^{-1}=\frac{1}{\lambda_{k}}=\frac{\overline{\lambda_{k}}}{\lambda_{k}\overline{\lambda_{k}}}=\frac{a_{k}-ib_{k}}{a_{k}^{2}+b_{k}^{2}}.$
where complex conjugate of a Gaussian integer $x+iy$ is denoted as $\overline{x+iy}=x-iy$.
Hence, $\lambda_{k}$ is simply a complex number with integer real $\operatorname{Re}(\cdot)$ an imaginary $\operatorname{Im}(\cdot)$ parts (i.e., a Gaussian integer, by definition): $\lambda_{k}\in\mathbb{Z}[i]$. In general, $\lambda_{k}^{-1}$ is not a Gaussian integer, but a Gaussian rational: $\lambda_{k}^{-1}\in\mathbb{Q}[i]$, which means that both real and imaginary parts of $\lambda_{k}^{-1}$ are rational numbers, or $\operatorname{Re}(\lambda_{k}^{-1}),\operatorname{Im}(\lambda_{k}^{-1})\in\mathbb{Q}$:
$\operatorname{Re}(\lambda_{k}^{-1})=\frac{a_{k}}{a_{k}^{2}+b_{k}^{2}},\>\operatorname{Im}(\lambda_{k}^{-1})=\frac{-b_{k}}{a_{k}^{2}+b_{k}^{2}}.$
The secret watermarking key $\lambda_{k}$, termed as challenge factor, could be fixed or change at each time interval, hence the subscript “kk” used in its definition. The changes on $\lambda_{k}$ can be scheduled according to the desired security level.
The embedding function, based on complex number arithmetic operations, is defined as follows. Let $\delta_{j,k}$ be the Gaussian integer $\delta_{j,k}=d_{j,k}+iw_{k}$.
Remark 3.
Note that $\delta_{j,k}$ is simply a complex number (Gaussian integer), whose real part is the original data and the imaginary part is the embedded watermark: $\operatorname{Re}(\delta_{j,k})=d_{j,k}$ and $\operatorname{Im}(\delta_{j,k})=w_{k}$.
As detailed in Definitions 1 and 2, both the data and the watermark are assumed to be integers and, hence, the complex number $\delta_{j,k}$ is indeed a Gaussian integer. $\triangle$
Now, the watermarked data [^1] We use the notation $\delta^{\prime}_{j,k}$ instead of $d^{\prime}_{j,k}$ to remark that the value is complex (Gaussian integer) and not a regular integer. $\delta^{\prime}_{j,k}$ can be obtained applying the following embedding function:
$\begin{split}E_{w}(d_{j,k},w_{k},\lambda_{k})&=\lambda_{k}\cdot\delta_{j,k}\\ &=(a_{k}+ib_{k})(d_{j,k}+iw_{k})\\ &=(a_{k}d_{j,k}-b_{k}w_{k})+i(a_{k}w_{k}+b_{k}d_{j,k}).\end{split}$ (1)
Remark 4.
$\delta^{\prime}_{j,k}$ is, again, a Gaussian integer, since it is the product of two Gaussian integers ($\lambda_{k}$ and $\delta_{j,k}$). Hence, the real and the imaginary parts of the watermarked data are integers. This is particularly interesting if we want to carry out further operations on this data, such as public-key encryption. $\triangle$
Remark 5.
From a purely mathematical point of view, there is no need to constrain the data, the watermark and the watermarking key to be integers or Gaussian integers. Consequently, a formulation of the scheme based on real and complex numbers (without constraints) is perfectly possible. However, in practical implementations, using floating point numbers would produce round-off errors in simple computations (such as multiplications or divisions). In addition, floating point numbers are not handled directly in some operations, such as public-key encryption. For these reasons, the proposed method is defined using integers in this paper.      $\triangle$
Remark 6.
It may be argued that the proposed approach is just a linear transformation of the vector $d_{j,k},w_{k}]{{}^{\mathrm{T}}}$ to a new vector $v$, obtained multiplying by a particular $2\times 2$ matrix, as follows:
$\left[\begin{array}[]{c}v_{1}\\ v_{2}\end{array}\right]=\left[\begin{array}[]{cc}a_{k}&-b_{k}\\ b_{k}&a_{k}\end{array}\right]=\left[\begin{array}[]{c}d_{j,k}\\ w_{k}\end{array}\right],$
where $v_{1}$ corresponds to $\operatorname{Re}(\delta^{\prime}_{j,k})$ and $v_{2}$ to $\operatorname{Im}(\delta^{\prime}_{j,k})$.
While this is true, the idea behind the proposed scheme is to exploit the algebraic operations that are already available for complex numbers. There are libraries for complex number operations in most programming languages and, hence, they can be directly applied in any practical implementation of the method. This prevents the explicit use of matrices and matrix inverses, which may be cumbersome. Obviously, the method can be further generalized in its vector form, as discussed in Section 3.4. $\triangle$
The extraction of the original data and the watermark from the watermarked data is straightforward, since the reversible extraction function $R_{w}$ can be defined as follows:
$\left[\begin{array}[]{c}d_{j,k}\\ w_{k}\end{array}\right]=\left[\begin{array}[]{c}\operatorname{Re}(\lambda_{k}^{-1}\delta^{\prime}_{j,k})\\ \operatorname{Im}(\lambda_{k}^{-1}\delta^{\prime}_{j,k})\end{array}\right].$ (2)
Consequently, we only need to multiply $\delta^{\prime}_{j,k}$ by $\lambda_{k}^{-1}$ (the complex inverse of $\lambda_{k}$) and, then, take the real and the imaginary parts to recover both the original data and the watermark.
Remark 7.
Expression (2) perfectly matches the formulation given for $R_{w}$ in Definition 3, since it takes, as input parameters, $\delta^{\prime}_{j,k}$ (analogous to $d^{\prime}_{j,k}$) and $K_{w}=\lambda_{k}$. $\triangle$
Remark 8.
The proposed method is vulnerable against Gaussian integer factorization. Even with just one eavesdropped value, a malicious observer may try to compute the Gaussian integer factors of the watermarked data and obtain knowledge about the original data, the watermark or the watermarking key. Consequently, for practical applications, it is not recommended to rely solely on the challenge $\lambda_{k}$ as a security measure. The scheme can be combined with random numbers and/or cryptographic functions for increased security and privacy, as shown in the protocols proposed in Section 5.
$\triangle$
Remark 9.
Although the data and the watermark are clearly separated in $\delta_{j,k}$ (the real part contains the data and the imaginary part is the watermark), this is no longer true after multiplying by $\lambda_{k}$. In the watermarked data $\delta^{\prime}_{j,k}=\lambda_{k}\cdot\delta_{j,k}$, the data $d_{j,k}$ and the watermark $w_{k}$ contribute to both the real and the imaginary parts. To recover the data and the watermark, knowledge about the secret key $\lambda_{k}$ is necessary (to compute $\lambda_{k}^{-1}$).
Hence, an eavesdropper who accessed $\delta^{\prime}_{j,k}$ would not be able to determine the data or the watermark directly, but may try to infer information from Gaussian integer factorization, as discussed in Remark 8. In addition, if the eavesdropper had access to the watermarked values corresponding to different time samples ($k$) or different devices ($j$), they could try to obtain even more information about $\lambda_{k}$, especially if it is not changed frequently.
$\triangle$
3.3 Data Aggregation
When a collection of devices $j=1,2,\dots,N$ transmit data, it is often desirable, for privacy reasons, to work on the aggregated data rather than individual readings. Let $d_{j,k}$ be the sensor data (an integer) from device $j$ at time $t_{k}$, and $S_{k}$ (Definition 4) be the total aggregated data.
For a data aggregation protocol, each device $j$ at time $t_{k}$ first combines the watermark $w_{k}$ with its data to form a complex number $\delta_{j,k}=d_{j,k}+iw_{k}$. This individual watermarked data is then scaled by the challenge factor $\lambda_{k}$ (which is provided by the Data Collector for the specific $k$-th round) to produce $\delta^{\prime}_{j,k}=\lambda_{k}\delta_{j,k}$.
The total aggregated watermarked data, denoted as $\sigma^{\prime}_{k}$, is the sum of these individual scaled and watermarked complex numbers across all devices:
$\sigma^{\prime}_{k}=\sum_{j=1}^{N}\delta^{\prime}_{j,k}=\sum_{j=1}^{N}\lambda_{k}\delta_{j,k}.$
Since $\lambda_{k}$ is a common factor for all devices in a given round, we can factor it out of the summation:
$\sigma^{\prime}_{k}=\lambda_{k}\sum_{j=1}^{N}\delta_{j,k}=\lambda_{k}\sum_{j=1}^{N}(d_{j,k}+iw_{k})=\lambda_{k}\sigma_{k},$
with $\sigma_{k}=\sum_{j=1}^{N}(d_{j,k}+iw_{k})$. By separating the real and imaginary components of the sum, and noting that the watermark $w_{k}$ is the same for all devices in a given round, the expression for $\sigma^{\prime}_{k}$ can be rearranged as:
$\sigma^{\prime}_{k}=\lambda_{k}\left(\sum_{j=1}^{N}d_{j,k}+i\sum_{j=1}^{N}w_{k}\right)=\lambda_{k}(S_{k}+iNw_{k}).$ (3)
where $S_{k}$ is the aggregated real data and $Nw_{k}$ is obtained as the aggregated imaginary part.
Similarly to the non-aggregated case, the result $\sigma^{\prime}_{k}$ is a complex number where both the aggregated data $S_{k}$ and the watermark $w_{k}$ (scaled by $N$) affect its real and imaginary parts. It is not possible to separate these values without the secret challenge factor $\lambda_{k}$.
The extraction of both the aggregated data $S_{k}$ and the watermark $w_{k}$ from $\sigma^{\prime}_{k}$ is straightforward. We can first multiply $\sigma^{\prime}_{k}$ by the inverse of $\lambda_{k}$ in the complex domain ($\lambda_{k}^{-1}$). Then, the real part of the result directly yields the aggregated original data ($S_{k}$), and the imaginary part, divided by the number of devices ($N$), gives the watermark ($w_{k}$). This extraction process can be formally expressed as:
$\left[\begin{array}[]{c}S_{k}\\ w_{k}\end{array}\right]=\left[\begin{array}[]{c}\operatorname{Re}(\lambda_{k}^{-1}\sigma^{\prime}_{k})\\ \displaystyle\frac{1}{N}\operatorname{Im}(\lambda_{k}^{-1}\sigma^{\prime}_{k})\end{array}\right].$ (4)
Hence, both the aggregated data and the watermark can be easily recovered from the aggregated-watermarked data.
A numerical example with small numbers is provided in Section 3.5 to illustrate how the embedding and recovery methods work, both for single data and for data aggregration.
Remark 10.
This particular version of the extraction function $R^{\mathrm{Agg}}_{w}$ does require $N$ as an input: $R^{\mathrm{Agg}}_{w}(\sigma^{\prime}_{k},\lambda_{k},N)$. However, if further constraints in the definition of the watermark and the number of devices are introduced, an alternative aggregated extraction function that would not require $N$ as an input may be obtained.
For example, if the number of devices $N$ was limited to a maximum of $100$: $N\leq 100$ (which may be enough in many applications), it would suffice to choose the watermark $w_{k}$ as any integer with no prime factors lower than 100. A partial factorization of $\operatorname{Im}(\lambda_{k}^{-1}\sigma^{\prime}_{k})=Nw_{k}$, for primes below 100, would clearly reveal both $N$ and $w_{k}$. All prime factors of $\operatorname{Im}(\lambda_{k}^{-1}\sigma^{\prime}_{k})$ lower than 100, with their respective exponents, would yield $N$. Once $N$ is known (by multiplying all those powers of small prime factors), then $w_{k}$ could be finally obtained by dividing $\operatorname{Im}(\lambda_{k}^{-1}\sigma^{\prime}_{k})=Nw_{k}$ by $N$.
In any case, knowledge about the number of participants is, in general, acceptable in many real-life situations. $\triangle$
3.4 Generalized Vector-based Approach for RDH
As discussed in Remark 6, a similar RDH scheme could be obtained using vectors and matrices instead of complex numbers. In this case, the complex number would be a particular example of the vector case (but with some advantages as already discussed).
In the generalized case, the watermarked data $d^{\prime}_{j,k}$ can be defined in vector form, as follows:
$d^{\prime}_{j,k}=\left[\begin{array}[]{cc}m_{11,k}&m_{12,k}\\ m_{21,k}&m_{22,k}\end{array}\right]\left[\begin{array}[]{c}d_{j,k}\\ w_{k}\end{array}\right]=\left[\begin{array}[]{c}m_{11,k}d_{j,k}+m_{12,k}w_{k}\\ m_{21,k}d_{j,k}+m_{22,k}w_{k}\end{array}\right],$
where a $2\times 2$ matrix makes the appropriate transformation.
For reversibility, it is enough that the matrix be invertible, i.e., that its determinant is different from zero: $m_{11,k}m_{22,k}-m_{12,k}m_{21,k}\neq 0$. If this condition is satisfied, the original data and the watermark can be extracted from $d^{\prime}_{j.k}$ as follows:
$\left[\begin{array}[]{c}d_{j,k}\\ w_{k}\end{array}\right]=\left[\begin{array}[]{cc}m_{11,k}&m_{12,k}\\ m_{21,k}&m_{22,k}\end{array}\right]^{-1}d^{\prime}_{j,k}.$
Since this is a linear transformation, it would also work for aggregated data, similarly as in the complex numbers case.
The property of “distributing” the watermark and the data in the two components of the resulting vector ($d^{\prime}_{j,k}$) will be satisfied as far as none of the elements of the matrix is zero. Again, if all the values $m_{11,k},\dots,m_{22,k}$ are integers, the obtained vector will also have integer components ($d^{\prime}_{j,k}\in\mathbb{Z}^{2}$), which may be convenient for different applications, especially those that require public-key encryption.
Remark 11.
This vector form of the reversible watermarking system, however, requires four parameters ($m_{11,k},\dots,m_{22,k}$) instead of two: the real ($a_{k}$) and imaginary ($b_{k}$) parts of a Gaussiann integer $\lambda_{k}$. While using four parameters offers no clear advantage over the two-parameter (complex number) case, it does require a larger secret watermarking key at both the embedding and extraction ends. This increased key size, without any compensating benefits, can be considered a disadvantage of the generalized approach, leading to the preference for the complex number-based formulation in this paper. The only advantage of such a scheme would rely on the longer apparent key size, but this can be easily replicated with two values instead of four just by selecting a larger range for each of them. $\triangle$
Obviously, the method could be further extended by introducing larger vectors and matrices. However, no clear advantages would be obtained with that additional complexity.
3.5 Numerical Example for Watermark Embedding, Extraction and Data Recovery
This section presents an example of how the method would work with simple data. In this example, three devices produce the following data: $d_{1,1}=5$, $d_{2,1}=8$, and $d_{3,1}=17$ at time $t_{1}$. All of them must embed the watermark $w_{1}=4$, and the secret embedding key is $\lambda_{1}=3+2i$. With this setting, the inverse $\lambda_{1}^{-1}$ can be obtained as follows:
$\lambda_{1}^{-1}=\frac{3-2i}{3^{2}+2^{2}}=\frac{3}{13}-\frac{2}{13}i.$
The corresponding embedded data are the following:
$\begin{split}\delta^{\prime}_{1,1}&=(3+2i)(5+4i)=7+22i,\\ \delta^{\prime}_{2,1}&=(3+2i)(8+4i)=16+28i,\\ \delta^{\prime}_{3,1}&=(3+2i)(17+4i)=43+46i.\end{split}$
Remark 12.
As discussed in Remark 8, a simple Gaussian integer factorization of eavesdropped data could reveal much information. For example, a factorization calculator for $7+22i$ yields $7+22i=(3+2i)(5+4i)$, which leaks $\lambda_{k}=3+2i$, in one factor, and the original data and watermark, in the other factor. While an individual eavesdropped data would not be enough to identify $\lambda_{k}$ and $d_{j,k}+iw_{k}$ from the computed factors, it is obvious that a small number of leaked data would suffice for an attacker. Additional security measures are required for the transmission of watermarked data. In addition, it is recommended to modify $\lambda_{k}$ frequently (e.g., at each time sample). $\triangle$
Remark 13.
As in the regular integer case (without complex numbers), the use of large prime factors for $\lambda_{k}$ would make the factorization problem harder for an attacker and, consequently, would increase security. $\triangle$
At the receiver end, the following computation would be performed:
$\begin{split}\lambda_{1}^{-1}\delta^{\prime}_{1,1}&=\left(\frac{3}{13}-\frac{2}{13}i\right)(7+22i)=5+4i,\\ \lambda_{1}^{-1}\delta^{\prime}_{2,1}&=\left(\frac{3}{13}-\frac{2}{13}i\right)(16+28i)=8+4i,\\ \lambda_{1}^{-1}\delta^{\prime}_{3,1}&=\left(\frac{3}{13}-\frac{2}{13}i\right)(43+46i)=17+4i.\end{split}$
Thus, taking the real and the imaginary parts after the multiplication recovers both the original data (5, 8 or 17) and the watermark (4), respectively.
Reversible extraction is also possible from aggregated data. The aggregated data for these three devices at time $t_{1}$ can be computed as:
$S_{1}=\sum_{j=1}^{3}d_{j,1}=d_{1,1}+d_{1,2}+d_{1,3}=5+8+17=30.$
Similarly, the aggregated watermarked data can be obtained as follows:
$\begin{split}\sigma^{\prime}_{1}&=\sum_{j=1}^{3}\delta^{\prime}_{j,1}=\delta^{\prime}_{1,1}+\delta^{\prime}_{1,2}+\delta^{\prime}_{1,3}\\ &=(7+22i)+(16+28i)+(43+46i)\\ &=66+96i.\end{split}$
The receiver end would compute the following:
$\lambda_{1}^{-1}\sigma^{\prime}_{1}=\left(\frac{3}{13}-\frac{2}{13}i\right)(66+96i)=30+12i,$
whose real part provides the aggregated data: $S_{1}=\operatorname{Re}(30+12i)=30$, whereas the imaginary part is $\operatorname{Im}(30+12i)=12=3\cdot 4=Nw_{1}$. If the imaginary part is divided by the number of devices ($N=3$), the embedded watermark ($w_{1}=4$) is finally extracted.
4 Homomorphic Encryption on the Complex Domain
This section presents the extension of ElGamal and Paillier cryptosystem to use complex-valued data.
The extension of ElGamal encryption to Gaussian integers requires a careful examination of the mathematical basis of the cryptosystem. The first building block of the standard ElGamal cryptosystem is the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$, for a prime number $p$, and a generator $g$ for such a group. In the Gaussian integer case, we must have a prime number $\pi$ in $\mathbb{Z}[i]$ and a consistent definition of the modular arithmetic $(\operatorname{\mathrm{mod}}\pi)$.
For a better understanding of the adaptation of this cryptosystem to the complex domain, a brief introduction to modular arithmetic over Gaussian integers is provided in A. A more detailed and comprehensive reference on Gaussian integers can be found in (Conrad, 2016).
4.1 ElGamal Encryption over Gaussian Integers
The principles of ElGamal encryption can be extended to other finite cyclic groups, including the multiplicative group of Gaussian integers modulo a Gaussian prime. This adaptation exploits the rich algebraic structure of Gaussian integers while maintaining the security reliance on the DLP.
We define the “Gaussian version” of ElGamal encryption as follows. Let $p$ be a Gaussian prime (preferably a regular prime integer of the form 4n14n-1, as discussed in A). Let $\gamma$ be a generator of the cyclic multiplicative group of Gaussian integers modulo $p$, denoted as ${\mathbb{Z}[i]}^{*}_{p}$. The ElGamal encryption scheme over Gaussian integers consists of the same three fundamental algorithms:
  • Key generation: The key generation process involves selecting a random integer $a$ such that $1\leq a\leq\mathcal{N}-1$ as the private key, where $\mathcal{N}$ is the order of the group ${\mathbb{Z}[i]}^{*}_{p}$. For a prime integer $p$ (of the form 4n14n-1), $\mathcal{N}=p^{2}-1$. Then, the public key $\mathcal{K}=\gamma^{a}\,(\operatorname{\mathrm{mod}}p)$ is computed. The public key is formed as the triple $\mathbf{K}=(p,\gamma,\mathcal{K})$.
  • Encryption: To encrypt a message $\mu\in{\mathbb{Z}[i]}^{*}_{p}$, a random integer $b$, such that $1\leq b\leq\mathcal{N}-1$, is selected as the ephemeral key. The ciphertext is then computed as $\psi_{1}=\gamma^{b}\,(\operatorname{\mathrm{mod}}p)$ and $\psi_{2}=\mu\cdot\mathcal{K}^{b}\,(\operatorname{\mathrm{mod}}p)$, resulting in the pair $(\psi_{1},\psi_{2})$. The encryption function can be denoted as follows:
    $(\psi_{1},\psi_{2})=\mathrm{Enc}_{\mathbf{K}}(\mu,b)=\left(\gamma^{b}\,(\operatorname{\mathrm{mod}}p),\mu\cdot\mathcal{K}^{b}\,(\operatorname{\mathrm{mod}}p)\right).$
  • Decryption: The decryption of the ciphertext $(\psi_{1},\psi_{2})$ using the private key $a$ involves computing the shared secret $\tau_{k}=\psi_{1}^{a}\,(\operatorname{\mathrm{mod}}p)$. Then, the message is decrypted as $\mu=\psi_{2}\cdot\tau_{k}^{-1}\,(\operatorname{\mathrm{mod}}p)$. This decryption function can be formally denoted as follows:
    $\mathrm{Dec}_{a}(\psi_{1},\psi_{2})=\psi_{2}\cdot(\psi_{1}^{a})^{-1}\,(\operatorname{\mathrm{mod}}p).$
The selection of a suitable generator $\gamma$ for the group ${\mathbb{Z}[i]}^{*}_{p}$ is a practical consideration, similar to the regular integer case. This can be efficiently achieved through probabilistic methods involving testing candidate elements.
The security of this Gaussian integer ElGamal variant fundamentally relies on the difficulty of the DLP in the group ${\mathbb{Z}[i]}^{*}_{p}$. If an attacker can efficiently solve the DLP in this group, they could compute the private key $a$ from the public key $\mathcal{K}$ and subsequently decrypt any ciphertext. For sufficiently large Gaussian primes $p$, the DLP is believed to be computationally infeasible.
As a key property, the order of this group, $\mathcal{N}=p^{2}-1$, is significantly larger than p1p-1 for the same numerical prime $p$ used in the regular integer ElGamal. This relationship, $\mathcal{N}=(p-1)(p+1)$, is particularly advantageous for security. If a prime $p$ is chosen such that p1p-1 possesses a large prime factor (a standard requirement for robust security in integer-based ElGamal against algorithms like Pohlig-Hellman (Pohlig and Hellman, 1978)), then this same large prime factor will inherently be a factor of $\mathcal{N}$. Therefore, the Gaussian integer ElGamal variant naturally inherits at least the same level of security (in terms of resistance to Pohlig-Hellman attacks) as its regular integer counterpart for a given $p$. Furthermore, the existence of the additional factor $(p+1)$ in $\mathcal{N}$ can potentially introduce even larger prime factors, further enhancing the security of the scheme. Thus, careful selection of the Gaussian prime $p$ to ensure that the largest prime factor of $\mathcal{N}$ is sufficiently large guides robust parameter choices.
Similar to its regular integer counterpart, this Gaussian integer ElGamal encryption exhibits a homomorphic property under component-wise multiplication of ciphertexts. This means that given two ciphertexts $\mathrm{Enc}_{\mathbf{K}}(\mu_{1},b_{1})=(\psi_{1},\psi_{2})$ and $\mathrm{Enc}_{\mathbf{K}}(\mu_{2},b_{2})=(\psi^{\prime}_{1},\psi^{\prime}_{2})$, multiplying them component-wise as $(\psi_{1}\cdot\psi^{\prime}_{1},\psi_{2}\cdot\psi^{\prime}_{2})$ results in a ciphertext that decrypts to the product of the original plaintexts:
$\mathrm{Dec}_{a}\left(\mathrm{Enc}_{\mathbf{K}}(\mu_{1},b_{1})\cdot\mathrm{Enc}_{\mathbf{K}}(\mu_{2},b_{2})\right)=\mu_{1}\cdot\mu_{2}\,(\operatorname{\mathrm{mod}}p).$
This property enables its use in various cryptographic protocols and specifically finds application in the proposed method for combining reversible watermarking and encryption in this paper.
4.1.1 Numerical example
This example demonstrates the ElGamal encryption process and its homomorphic property using Gaussian integers. The multiplicative group of Gaussian integers modulo $p=23$, denoted as ${\mathbb{Z}[i]}^{*}_{p}$, is selected. The order of this group is $\mathcal{N}=23^{2}-1=528$.
For the parameters, the prime modulus $p=23$ and the generator $\gamma=1+2i$ are taken, and the private key is chosen as $a=7$. Now, $\mathcal{K}$ is computed as
$\mathcal{K}=\gamma^{a}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=6+2i,$
and the public key is thus formed as the triple $\mathbf{K}=(p,\gamma,\mathcal{K})=(23,1+2i,6+2i)$.
To encrypt a first message $\mu_{1}=5+4i$, an ephemeral key $b_{1}=5$ is chosen. The ciphertext components are computed as:
$\begin{split}\psi_{1}&=\gamma^{b_{1}}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{5}\,(\operatorname{\mathrm{mod}}23)=18+8i,\\ \psi_{2}&=\mu_{1}\cdot\mathcal{K}^{b_{1}}\,(\operatorname{\mathrm{mod}}p)=(5+4i)\cdot(6+2i)^{5}\,(\operatorname{\mathrm{mod}}23)=21+11i.\end{split}$
The resulting ciphertext is $(\psi_{1},\psi_{2})=(18+8i,21+11i)$. Decryption of this ciphertext using the private key $a=7$ successfully yields $\mu_{1}=(5+4i)\,(\operatorname{\mathrm{mod}}23)$.
For a second message $\mu_{2}=3+2i$, an ephemeral key $b_{2}=7$ is chosen. The ciphertext components are obtained as:
$\begin{split}\psi^{\prime}_{1}&=\gamma^{b_{2}}\,(\operatorname{\mathrm{mod}}p)=(1+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=6+2i,\\ \psi^{\prime}_{2}&=\mu_{2}\cdot\mathcal{K}^{b_{2}}\,(\operatorname{\mathrm{mod}}p)=(3+2i)\cdot(6+2i)^{7}\,(\operatorname{\mathrm{mod}}23)=21+16i.\end{split}$
The resulting ciphertext is $(\psi^{\prime}_{1},\psi^{\prime}_{2})=(6+2i,21+16i)$. Decryption of this ciphertext using the private key $a=7$ successfully yields $\mu_{2}=(3+2i)\,(\operatorname{\mathrm{mod}}23)$.
The homomorphic property is demonstrated by multiplying the two ciphertexts component-wise: $(\psi_{1},\psi_{2})=(18+8i,21+11i)$ and $(\psi^{\prime}_{1},\psi^{\prime}_{2})=(6+2i,21+16i)$. The resulting ciphertext components, obtained by component-wise multiplication, are:
$\begin{split}\psi^{\prime\prime}_{1}&=\psi_{1}\cdot\psi^{\prime}_{1}=(18+8i)\cdot(6+2i)=(0+15i)\,(\operatorname{\mathrm{mod}}23),\\ \psi^{\prime\prime}_{2}&=\psi_{2}\cdot\psi^{\prime}_{2}=(21+11i)\cdot(21+16i)=(12+15i)\,(\operatorname{\mathrm{mod}}23).\end{split}$
The combined ciphertext for the product is then expressed as $(\psi^{\prime\prime}_{1},\psi^{\prime\prime}_{2})=(0+15i,12+15i)$.
Decryption of this combined ciphertext $(\psi^{\prime\prime}_{1},\psi^{\prime\prime}_{2})=(0+15i,12+15i)$ using the private key $a=7$ yields $\mu^{\prime\prime}=(7+22i)\,(\operatorname{\mathrm{mod}}23)$. This precisely matches the direct multiplication of the original plaintexts:
$\begin{split}\mu_{1}\cdot\mu_{2}&=(5+4i)\cdot(3+2i)\,(\operatorname{\mathrm{mod}}23)\\ &=\left((15-8)+(10+12)i\right)\,(\operatorname{\mathrm{mod}}23)\\ &=(7+22i)\,(\operatorname{\mathrm{mod}}23).\end{split}$
This confirms the correct operation and homomorphic property of ElGamal encryption over Gaussian integers.
4.2 Component-wise Paillier Encryption of Gaussian Integers
While the previously discussed ElGamal scheme was extended to operate natively over Gaussian integers, the additive homomorphic property of Paillier encryption allows for a straightforward application to complex numbers (specifically, Gaussian integers) without requiring a specialized “complex version” of the cryptosystem. This is achieved by encrypting the real and imaginary parts of a Gaussian integer independently.
Let a Gaussian integer plaintext message be represented as $\mu=m_{R}+im_{I}$, where $m_{R}$ and $m_{I}$ are integers. To encrypt $\mu$ using the standard Paillier scheme with public key $\mathsf{K_{pub}}=(n,g)$, the process is as follows:
  • Each component is encrypted separately:
    $c_{R}$ $=\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{R},r_{R})=g^{m_{R}}\cdot r_{R}^{n}\,(\operatorname{\mathrm{mod}}n^{2}),$
    $c_{I}$ $=\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{I},r_{I})=g^{m_{I}}\cdot r_{I}^{n}\,(\operatorname{\mathrm{mod}}n^{2}),$
    where $r_{R}$ and $r_{I}$ are randomly chosen integers for each encryption, satisfying $0<r_{R},r_{I}<n$ and $\mathrm{gcd}(r_{R},n)=\mathrm{gcd}(r_{I},n)=1$.
  • The ciphertext for the Gaussian integer $\mu$ is then represented as a pair of Paillier ciphertexts: $(c_{R},c_{I})$.
Decryption of a Gaussian integer ciphertext $(c_{R},c_{I})$ is performed component-wise using the Paillier private key $\mathsf{K_{priv}}=(L_{P},M_{P})$: $m_{R}=\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{R})$ and $m_{I}=\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{I})$. The decrypted Gaussian integer is then reconstructed as $\mu=m_{R}+im_{I}$.
The primary advantage of this component-wise approach is that it directly inherits the additive homomorphic property of the standard Paillier scheme. If we have two encrypted Gaussian integer messages, $\mathrm{Enc}(\mu_{1})=(c_{1R},c_{1I})$ for $\mu_{1}=m_{1R}+im_{1I}$ and $\mathrm{Enc}(\mu_{2})=(c_{2R},c_{2I})$ for $\mu_{2}=m_{2R}+im_{2I}$, their sum can be computed in the encrypted domain as follows:
  • For the real part: $c_{\mathrm{sum},R}=c_{1R}\cdot c_{2R}\,(\operatorname{\mathrm{mod}}n^{2})$.
  • For the imaginary part: $c_{\mathrm{sum},I}=c_{1I}\cdot c_{2I}\,(\operatorname{\mathrm{mod}}n^{2})$.
When the resulting ciphertext $(c_{\mathrm{sum},R},c_{\mathrm{sum},I})$ is decrypted, the real and imaginary components sum independently:
$\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{\mathrm{sum},R})$ $=(m_{1R}+m_{2R})\,(\operatorname{\mathrm{mod}}n),$
$\mathrm{Dec}_{\mathsf{K_{priv}}}(c_{\mathrm{sum},I})$ $=(m_{1I}+m_{2I})\,(\operatorname{\mathrm{mod}}n).$
Thus, the decryption yields the sum of the original Gaussian integers: $\mu_{1}+\mu_{2}=(m_{1R}+m_{2R})+i(m_{1I}+m_{2I})\,(\operatorname{\mathrm{mod}}n)$.
This component-wise additive homomorphism is useful for applications such as privacy-preserving aggregation of complex-valued data, where sums of watermarked sensor readings can be computed without revealing individual values. The security of this approach remains equivalent to that of the underlying regular Paillier encryption scheme.
5 Protocols for Joint Watermarking and Encryption on the Complex Domain
This section presents the Hiding in the Imaginary Domain with Data Encryption (H[i][i]dden) framework, a novel approach designed to simultaneously embed information (watermarks) into data and encrypt it. Within this general framework, we detail two distinct protocols tailored for different application scenarios and taking advantage of different homomorphic properties.
The first protocol, termed H[i][i]dden-EG, uses Gaussian ElGamal encryption to provide joint watermarking and encryption for individual data points. The second one, denoted as the H[i][i]dden-AggP protocol, uses Paillier encryption’s additive homomorphism to enable joint watermarking, encryption, and privacy-preserving aggregation of data from multiple sources. Each protocol is presented with its initialization and data exchange phases, and also their respective complexities.
5.1 System Model and Security Requirements
Establishing a rigorous security foundation for dispersed data in distributed environments requires a clear definition of the participating entities and the adversarial assumptions. The H[i][i]dden framework is specifically designed to address the vulnerabilities inherent in resource-constrained networks, where data must be transmitted across potentially insecure channels while maintaining perfect reversibility. In this section, we formalize the system model as a precursor to the detailed analysis of specific threat vectors and protocol-specific defenses.
5.1.1 System Entities
The H[i][i]dden framework involves the following primary entities:
  • Sensors ($\mathrm{S}_{j}$): Distributed, resource-constrained devices that generate original scalar data $d_{j,k}$. They perform the reversible embedding of the watermark $w_{k}$ and the initial encryption steps.
  • Partial Aggregators ($\mathrm{S}_{\ell}$): In the H[i][i]dden-AggP (Paillier-based) version, certain sensors or intermediate nodes act as aggregators. They use homomorphic properties to sum ciphertexts without accessing plaintext data.
  • Data Collector (DC): The central entity that initializes the system parameters, distributes challenge factors $\lambda_{k}$, and performs the final decryption to recover both the fine-grained (or aggregated) data and the watermark.
5.1.2 Security and Privacy Requirements
The following properties define the core security goals of the framework.
Definition 6 (Confidentiality).
The sensor data $d_{j,k}$ and challenge factors $\lambda_{k}$ must remain protected from unauthorized parties. This is ensured by the semantic security of the Gaussian ElGamal and Paillier encryption schemes.
Definition 7 (Integrity and Accuracy).
The system must ensure that data has not been altered during transmission or aggregation. Accuracy is guaranteed by the perfect reversibility of the complex-domain embedding, while integrity is verified through the successful extraction of the pre-agreed watermark $w_{k}$.
Definition 8 (Data Provenance).
The framework must provide proof of the origin of the data. Since the watermark $w_{k}$ is secretly agreed upon between the DC and the legitimate sensors $\mathrm{S}_{j}$, the extraction of the correct watermark serves as a cryptographic proof that the data originated from an authorized source.
Definition 9 (Privacy).
Privacy is defined as the inability of any entity (other than the DC) to learn individual sensor readings. This is maintained even during aggregation via additive homomorphic encryption.
5.1.3 Attack Model
To analyze the security of the framework, we establish the following assumptions and adversarial capabilities:
  1. a)
    Entity Assumptions: The sensors $\mathrm{S}_{j}$ are assumed to be honest-but-curious; they follow the protocol but may attempt to learn information about other sensors’ data. The Data Collertor (DC) is assumed to be a trusted entity.
  2. b)
    Adversary Capabilities: We assume a potential Attacker ($\mathcal{A}$) that can be either passive or active. $\mathcal{A}$ may reside on any communication line in the system.
  3. c)
    Threat Vectors: Within this model, the system must defend against:
    • Eavesdropping: Interception of ciphertexts to compromise confidentiality.
    • Man-in-the-Middle (MitM): Active modification of messages between entities.
    • Masquerading: Attempting to impersonate a legitimate sensor.
    • Replay Attacks: Re-sending valid messages from previous rounds to skew current data.
    • False Data Injection (FDI): Injecting malicious readings into the individual or aggregated stream.
Figure 1: Graphical representation of the H[i][i]dden-EG protocol, including the initialization (above the horizontal dashed line) and the data exchange (below the horizontal dashed line) phases.
5.2 H[i][i]dden-EG: Joint Watermarking and Encryption for Individual Data
This section describes a protocol for secure and reversible information hiding between a sensor (S) and the Data Collector (DC). It applies the proposed complex domain framework and Gaussian ElGamal encryption to ensure sensor data ($d_{k}$) is transmitted confidentially while simultaneously embedding a time-dependent watermark ($w_{k}$) for verification and integrity.
Remark 14.
Although the data $d_{j,k}$ and the watermark $w_{k}$ were assumed to be simply integers (Definitions 1 and 2), the constraint of working modulo $p$ for a Gaussian prime satisfying $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$ in the H[i][i]dden-EG protocol imposes further constraints on the ranges of the data, the watermark, and the challenge factor. If both positive and negative values are allowed, the Gaussian integers $(\operatorname{\mathrm{mod}}p)$ can represent any value in:
${\mathbb{Z}[i]}^{*}_{p}=\left\{x+iy\mid x,y\in\left\{-\frac{p-1}{2},-\frac{p-1}{2}+1,\dots,\frac{p-1}{2}\right\}\text{ and }(x,y)\neq(0,0)\right\}.$
Hence, the following condition is required to avoid wrap-around: $\max(|\operatorname{Re}(\lambda_{k}\delta_{k})|,|\operatorname{Im}(\lambda_{k}\delta_{k})|)<p/2.$ $\triangle$
The protocol proceeds in two distinct phases: Initialization and watermarking and data exchange. The initialization phase establishes the necessary cryptographic keys and shared secrets between the DC and S:
  1. 1.
    Public Key Distribution: DC generates a Gaussian ElGamal key pair. It computes and publicly distributes its public key $\mathbf{K}=(p,\gamma,\mathcal{K})$.,
  2. 2.
    Watermark Agreement: DC and S agree on a set of shared watermarking keys or a mechanism to generate sequential, time-dependent watermarks: $w_{1},w_{2},\ldots,w_{M}$ corresponding to times $t_{1},t_{2},\ldots,t_{M}$. Each $w_{j}$ is an integer represented with $B$ bits (e.g., taking values in the range $0,\ldots,2^{B}-1$ for unsigned watermarks, or $2^{B-1},-2^{B-1}+1,\dots,2^{B-1}-1$ for signed ones).
Then, for the watermarking and data exchange phase, at each designated time $t_{k}$, sensor data is generated, watermarked, encrypted, and transmitted for verification.
  1. 1.
    DC’s Challenge Factor: DC chooses a Gaussian integer factor $\lambda_{k}$ with $\operatorname{Re}(\lambda_{k})\cdot\operatorname{Im}(\lambda_{k})\neq 0$ (from ${\mathbb{Z}[i]}^{*}_{p}$). DC then encrypts $\lambda_{k}$ using its own public key $\mathbf{K}$, yielding the ciphertext $(\psi_{k,1},\psi_{k,2})=\mathrm{Enc}_{\mathbf{K}}(\lambda_{k},b_{\lambda_{k}})$, where $b_{\lambda_{k}}$ is a randomly chosen ephemeral key. DC sends this ciphertext to S.
  2. 2.
    Sensor’s Data and Watermark Embedding: S generates its data reading $d_{k}$ (an integer, typically representing a real-world measurement). S then assigns the pre-agreed watermark $w_{k}$ to the imaginary part of the data, forming a complex-valued plaintext $\delta_{k}=d_{k}+iw_{k}$. Next, S encrypts this combined plaintext $\delta_{k}$ using DC’s public key $\mathbf{K}$, resulting in the ciphertext $(\psi^{\prime}_{k,1},\psi^{\prime}_{k,2})=\mathrm{Enc}_{\mathbf{K}}(\delta_{k},b_{\delta_{k}})$, where $b_{\delta_{k}}$ is another randomly chosen ephemeral key. Making use of the homomorphic property of Gaussian ElGamal, S component-wise multiplies the received challenge ciphertext with its own generated ciphertext:
    $\psi^{\prime\prime}_{k,1}$ $=\psi_{k,1}\cdot\psi^{\prime}_{k,1}\,(\operatorname{\mathrm{mod}}p),$
    $\psi^{\prime\prime}_{k,2}$ $=\psi_{k,2}\cdot\psi^{\prime}_{k,2}\,(\operatorname{\mathrm{mod}}p).$
    This step simultaneously performs both embedding and encryption, with the additional benefit of homomorphic multiplication. S then sends the resulting combined ciphertext $(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$ to the DC.
  3. 3.
    DC’s Decryption and Verification: Upon receiving the combined ciphertext $(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$, the DC decrypts it using its private key $a$. Due to the homomorphic property, the decrypted value $\mu_{k}=\mathrm{Dec}_{a}(\psi^{\prime\prime}_{k,1},\psi^{\prime\prime}_{k,2})$ will correspond to the product of the two original plaintexts: $\mu_{k}=\lambda_{k}\cdot\delta_{k}\,(\operatorname{\mathrm{mod}}p)$. Given the careful selection of $p$ to be sufficiently large, this modular result $\mu_{k}$ is guaranteed to be equivalent to the exact, non-modular product $\lambda_{k}\cdot\delta_{k}$ as Gaussian integers. To isolate the watermarked data, DC then computes $\delta_{k}=\mu_{k}/\lambda_{k}$ using standard complex number division. Finally, DC extracts the original sensor data $d_{k}$ from the real part of $\delta_{k}$ and the embedded watermark $w_{k}$ from its imaginary part. The data $d_{k}$ is accepted only if the extracted watermark $w_{k}$ correctly matches the expected time-dependent watermark; otherwise, the data is discarded, indicating potential tampering or an unauthorized source.
A graphical representation of the protocol is shown in Fig. 1.
5.2.1 Complexity analysis
Regarding computational complexity, the number of complex modular exponentiations and integer modular exponentiations, required at each party during the data exchange phase, is analyzed below.
For a Gaussian integer $\chi=x+iy\in{\mathbb{Z}[i]}^{*}_{p}$, its modular inverse $\chi^{-1}\,(\operatorname{\mathrm{mod}}p)$ is typically computed as $\overline{\chi}\cdot(\mathrm{N}(\chi))^{-1}\,(\operatorname{\mathrm{mod}}p)$. This involves obtaining the modular inverse of the integer norm $\mathrm{N}(\chi)=x^{2}+y^{2}$ modulo $p$ (an integer modular exponentiation), followed by a complex multiplication (Fermat’s Little Theorem):
$(x^{2}+y^{2})^{-1}\,(\operatorname{\mathrm{mod}}p)=(x^{2}+y^{2})^{p-2}\,(\operatorname{\mathrm{mod}}p).$
At the Sensor: In step 2 of the data exchange phase, S computes $\psi^{\prime}_{k,1}=\gamma^{b_{\delta_{k}}}\,(\operatorname{\mathrm{mod}}p)$ and $\mathcal{K}^{b_{\delta_{k}}}\,(\operatorname{\mathrm{mod}}p)$ (for $\psi^{\prime}_{k,2}$). Both of these are complex modular exponentiations. Therefore, the Sensor performs 2 complex modular exponentiations per data exchange.
At the Data Collector: In step 1, DC computes $\gamma^{b_{\lambda_{k}}}\,(\operatorname{\mathrm{mod}}p)$ and $\mathcal{K}^{b_{\lambda_{k}}}\,(\operatorname{\mathrm{mod}}p)$ for the challenge factor’s encryption. This amounts to 2 complex modular exponentiations. In step 3, DC performs the decryption and subsequent isolation of $\delta_{k}$:
  • Decryption: This involves computing $(\psi^{\prime\prime}_{k,1})^{a}\,(\operatorname{\mathrm{mod}}p)$, which is 1 complex modular exponentiation.
  • Computation of $\tau_{k}^{-1}$: To compute the inverse of the shared secret $\tau_{k}=(\psi^{\prime\prime}_{k,1})^{a}$, it is calculated as $\overline{\tau_{k}}\cdot(\mathrm{N}(\tau_{k}))^{-1}\,(\operatorname{\mathrm{mod}}p)$. This involves 1 integer modular exponentiation (for $(\mathrm{N}(\tau_{k}))^{-1}\,(\operatorname{\mathrm{mod}}p)$).
  • Computation of $\mu_{k}$: This is $\psi^{\prime\prime}_{k,2}\cdot\tau_{k}^{-1}\,(\operatorname{\mathrm{mod}}p)$, involving complex multiplication.
  • Computation of $\delta_{k}$: As $\mu_{k}$ is now the exact product $\lambda_{k}\cdot\delta_{k}$, DC computes $\delta_{k}=\mu_{k}/\lambda_{k}$ using standard complex number division.
In summary, per data exchange, the Data Collector performs a total of 3 complex modular exponentiations (2 for encrypting $\lambda_{k}$ and 1 for computing $(\psi^{\prime\prime}_{k,1})^{a}$) and 1 integer modular exponentiation (for the inverse of $\mathrm{N}(\tau_{k})$).
Remark 15.
Regarding the relative computational cost, a complex modular exponentiation is approximately four times as expensive as an integer modular exponentiation of a similar exponent length. This is primarily because each complex modular multiplication, which forms the building block of exponentiation, involves four integer modular multiplications:
$(a+ib)(c+id)\,(\operatorname{\mathrm{mod}}p)=((ac-bd)+i(ad+bc))\,(\operatorname{\mathrm{mod}}p).$
$\triangle$
The computational cost for each entity during a single data exchange round is analyzed in terms of equivalent integer modular exponentiations (as defined in Remark 15). The Sensor performs approximately 8 equivalent integer modular exponentiations for encrypting its watermarked data. In contrast, the Data Collector incurs a higher computational burden, totaling approximately 13 equivalent integer modular exponentiations per round (and sensor). The majority of this computational load at the DC is primarily due to the operations involved in encrypting the challenge factor ($\lambda_{k}$) sent to the Sensor, decrypting the combined ciphertext received from the Sensor, and then isolating the original watermarked data. This distribution of workload, where the Sensor’s load is comparatively lighter, aligns well with typical sensor network architectures where sensors are often resource-constrained devices, while the data collector usually possesses greater computational power.
Figure 2: Graphical representation of the H[i][i]dden-AggP protocol, including the initialization (above the horizontal dashed line) and the data exchange (below the horizontal dashed line) phases.
5.3 H[i][i]dden-AggP: Joint Watermarking, Encryption, and Aggregation
This section describes a protocol enabling privacy-preserving data aggregation from multiple sensors ($\mathrm{S}_{j}$, $j=1,\dots,N$) by a Data Collector (DC), incorporating reversible information hiding and applying Paillier encryption for additive homomorphism. The protocol ensures that aggregated sensor data is obtained securely, while a time-dependent watermark allows for verification and integrity checking.
The initialization phase establishes the cryptographic keys and shared secrets required for secure communication and data aggregation.
1) Paillier Public Key Distribution: DC generates a Paillier key pair. It computes and publicly distributes its public key $\mathsf{K_{pub}}=(n,g)$, where $n=pq$ for large primes $p,q$, and $g$ is a generator as defined in Section 2.2.2. DC keeps its corresponding private key $\mathsf{K_{priv}}=(L_{P},M_{P})$ secret.
2) Symmetric Key Exchange (for DC to Sensor): Each sensor $\mathrm{S}_{j}$ ($j=1,\ldots,N$), $\mathrm{S}_{j}$ generates a unique, random symmetric key $K_{\mathrm{sym},j}$ (e.g., an AES key). $\mathrm{S}_{j}$ then encrypts $K_{\mathrm{sym},j}$ using DC’s public Paillier key $\mathsf{K_{pub}}$, resulting in ciphertext $c_{j}=\mathrm{Enc}_{\mathsf{K_{pub}}}(K_{\mathrm{sym},j},r_{j})$, where $r_{j}$ is a random integer. $\mathrm{S}_{j}$ transmits $c_{j}$ to DC. Upon reception, DC decrypts $c_{j}$ using its private key $\mathsf{K_{priv}}$ to obtain $K_{\mathrm{sym},j}$. This process establishes a unique symmetric key $K_{\mathrm{sym},j}$ shared securely between DC and each individual sensor $\mathrm{S}_{j}$. These symmetric keys will subsequently be used for secure communication from DC to $\mathrm{S}_{j}$.
3) Watermark Agreement: DC and all sensors $\mathrm{S}_{j}$ collectively agree on a set of shared, sequential, time-dependent watermarks: $w_{1},w_{2},\ldots,w_{M}$ corresponding to times $t_{1},t_{2},\ldots,t_{M}$. Each $w_{k}$ (for the $k$-th round) is an integer represented with $B$ bits, which may be signed or unsigned as detailed in the H[i][i]dden-EG case.
At each designated time $t_{k}$, sensor data is generated, watermarked, encrypted, and aggregated for privacy-preserving transmission and verification:
1) DC’s Challenge Factor Distribution: DC chooses a Gaussian integer factor $\lambda_{k}$ with $\operatorname{Re}(\lambda_{k})\cdot\operatorname{Im}(\lambda_{k})\neq 0$. The real and imaginary parts of $\lambda_{k}$ must be chosen such that the resulting components of $\delta^{\prime}_{j,k}$ can be represented within Paillier’s plaintext modulus $n$. For each sensor $\mathrm{S}_{j}$ ($j=1,\ldots,N$), DC encrypts $\lambda_{k}$ using the pre-established unique symmetric key $K_{\mathrm{sym},j}$ shared with that specific sensor. Let this symmetric encryption be $\mathrm{Enc}_{K_{\mathrm{sym},j}}(\lambda_{k})$ (implying $\lambda_{k}$ is suitably serialized for symmetric encryption). DC then sends $\mathrm{Enc}_{K_{\mathrm{sym},j}}(\lambda_{k})$ to each $\mathrm{S}_{j}$. Each sensor $\mathrm{S}_{j}$ decrypts the received message using its $K_{\mathrm{sym},j}$ to obtain the plaintext $\lambda_{k}$.
2) Sensor Data Embedding and Initial Paillier Encryption: Each sensor $\mathrm{S}_{j}$ generates its data reading $d_{j,k}$ (an integer). $\mathrm{S}_{j}$ then combines the pre-agreed watermark $w_{k}$ with its data, forming a complex number $\delta_{j,k}=d_{j,k}+iw_{k}$. Using the received $\lambda_{k}$ (once decrypted), $\mathrm{S}_{j}$ then computes the scaled and watermarked complex number $\delta^{\prime}_{j,k}=\lambda_{k}\delta_{j,k}$. This $\delta^{\prime}_{j,k}$ is the value that will be aggregated. $\mathrm{S}_{j}$ encrypts $\delta^{\prime}_{j,k}$ component-wise using DC’s public Paillier key $\mathsf{K_{pub}}$. Let $\delta^{\prime}_{j,k}=m_{j,k,R}+im_{j,k,I}$ (where $m_{j,k,R}$ and $m_{j,k,I}$ are integers). The ciphertext is $(c_{j,k,R},c_{j,k,I})=(\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{j,k,R},r_{j,k,R}),\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{j,k,I},r_{j,k,I}))$, where $r_{j,k,R}$ and $r_{j,k,I}$ are fresh random numbers for each encryption.
3) Cascading Encrypted Summation (Tree Aggregation): The individual encrypted watermarked data points are then aggregated in a cascading manner, typically following a binary tree structure involving $\lceil\log_{2}(N)\rceil$ steps:
  • Initial Layer: Each $\mathrm{S}_{j}$ that is a leaf node in the aggregation tree generates its ciphertext $(c_{j,k,R},c_{j,k,I})$ as described in step 2.
  • Aggregation Steps: For each internal node in the aggregation tree, a designated sensor (or an intermediate aggregator) receives two ciphertexts, say $(C_{1R},C_{1I})=({c_{j_{1},k,R},c_{j_{1},k,I})}$ and $(C_{2R},C_{2I})=({c_{j_{2},k,R},c_{j_{2},k,I})}$, from its children (or from other sensors). Taking advantage of the additive homomorphic property of Paillier encryption, this aggregator computes the sum of the encrypted real parts ($C_{\mathrm{sum},R}=C_{1R}\cdot C_{2R}\,(\operatorname{\mathrm{mod}}n^{2})$) and the sum of the encrypted imaginary parts ($C_{\mathrm{sum},I}=C_{1I}\cdot C_{2I}\,(\operatorname{\mathrm{mod}}n^{2})$). The resulting ciphertext $(C_{\mathrm{sum},R},C_{\mathrm{sum},I})$ is then passed up the tree. This process continues until a single final encrypted sum is obtained.
  • Handling $N\neq 2^{X}$: If the number of sensors $N$ is not a power of 2, some “blank” or identity ciphertexts (e.g., encryption of $0+i0$) may be included in the aggregation process to complete the tree structure, or the tree may be handled as an unbalanced structure.
The final encrypted sum, $\sigma^{\prime}_{\mathrm{sum},k}=(C_{\mathrm{final},R},C_{\mathrm{final},I})$, is obtained by the root sensor in the aggregation tree (e.g., $\mathrm{S}_{1}$ if sensors are ordered by aggregation path) and is then sent to DC.
4) DC’s Decryption and Verification: Upon receiving the final aggregated ciphertext $\sigma^{\prime}_{\mathrm{sum},k}=(C_{\mathrm{final},R},C_{\mathrm{final},I})$, DC decrypts both components using its private Paillier key $\mathsf{K_{priv}}$. This yields the real and imaginary parts of the aggregated watermarked data:
$\sigma^{\prime}_{k,R}$ $=\mathrm{Dec}_{\mathsf{K_{priv}}}(C_{\mathrm{final},R})=\sum_{j=1}^{N}m_{j,k,R}\,(\operatorname{\mathrm{mod}}n),$
$\sigma^{\prime}_{k,I}$ $=\mathrm{Dec}_{\mathsf{K_{priv}}}(C_{\mathrm{final},I})=\sum_{j=1}^{N}m_{j,k,I}\,(\operatorname{\mathrm{mod}}n).$
The decrypted value is the aggregated watermarked data $\sigma^{\prime}_{k}=\sigma^{\prime}_{k,R}+i\sigma^{\prime}_{k,I}$. As derived in Section 3.3, $\sigma^{\prime}_{k}=\lambda_{k}(S_{k}+iNw_{k})$, where $S_{k}=\sum_{j=1}^{N}d_{j,k}$. DC then isolates the true aggregated data $S_{k}$ and the watermark $w_{k}$ by computing $(S_{k}+iNw_{k})=\sigma^{\prime}_{k}/\lambda_{k}$ using standard complex number division. Finally, DC extracts $S_{k}$ from the real part and obtains $w_{k}$ by dividing the imaginary part by $N$. If the extracted watermark $w_{k}$ correctly matches the expected time-dependent watermark, the aggregated data $S_{k}$ is accepted; otherwise, it is discarded, indicating potential tampering or an unauthorized source.
Remark 16.
The component-wise application of Paillier encryption requires the real and imaginary parts of the aggregate $\mu=\lambda_{k}(S_{k}+iNw_{k})$ to be handled modulo $n$. In this scenario, the additive homomorphic property allows for the accumulation of $N$ sensor readings. To ensure that the aggregate result does not exceed the capacity of the modulus $n$ and avoids modular wrap-around, the parameters must satisfy:
$\max(|\operatorname{Re}(\mu)|,|\operatorname{Im}(\mu)|)<n/2.$
This condition ensures that both components of the aggregate Gaussian integer can be correctly mapped back from the centered representative system of $\mathbb{Z}_{n}$ to their exact integer values in $\mathbb{Z}$, which is a prerequisite for the successful application of the perfect reversibility proven in Theorem 6. $\triangle$
Figure 3: Cascading summation protocol: example for 11 sensors.
A graphical representation of the H[i][i]dden-AggP protocol is shown in Fig. 2, whereas Fig. 3 illustrates how the cascading summation would work for 11 sensors. In this figure, the $\mathrm{S}_{1}$ sensor would obtain the final aggregate result to send it to the DC. It must be taken into account that the $\Sigma$” symbols in the figure represent multiplication of ciphertexts for the Paillier cryptosystem.
Remark 17.
The cascading aggregation protocol shown in Fig. 3 offers the advantage of performing the necessary summation internally within the sensor network without requiring a dedicated external entity. While this approach distributes the computational load, the topology implies a sequential delay of $\lceil\log_{2}(N)\rceil$ communication steps as partial aggregates propagate. Alternatively, if a participating sensor (e.g., $\mathrm{S}_{1}$) is sufficiently equipped with storage and computational power, it can act as a hub, receiving all ciphertexts directly from other nodes. While this reduces the transmission latency to a single step, the node must still perform N1N-1 modular multiplications (modulo $n^{2}$) sequentially to produce the final result. In fact, whether using a tree or a star arrangement, the algebraic work remains consistent. Performing aggregation within the sensor network itself rather than employing an external Data Aggregator (DA) minimizes the attack surface and circumvents the complex security considerations of third-party collusion (e.g., potential collusions between the DA and the DC). $\triangle$
5.3.1 Complexity Analysis
This section analyzes the computational and communication complexity of the proposed data aggregation protocol. We focus on the operations within a single data exchange round ($t_{k}$), as initialization costs are typically one-time or amortized over many rounds.
For computational complexity. the dominant costs arise from Paillier encryption and decryption operations, which primarily involve modular exponentiations modulo $n^{2}$. Other operations, such as symmetric encryption/decryption, complex modular multiplications (including those from homomorphic ciphertext additions), and standard complex divisions, are considered computationally negligible in this high-level analysis.
At the Sensor ($\mathrm{S}_{j}$): Each sensor $\mathrm{S}_{j}$ performs one initial encryption of its watermarked data $\delta^{\prime}_{j,k}$. Since this involves encrypting both the real and imaginary components separately using Paillier encryption, each sensor performs 2 Paillier encryptions per round. Each Paillier encryption typically involves two modular exponentiations (one for $g^{m}$ and one for $r^{n}$), totaling 4 modular exponentiations modulo $n^{2}$ per sensor for initial encryption.
If a sensor also acts as an intermediate aggregator in the tree, it performs additional Paillier ciphertext multiplications (homomorphic additions). While these are critical for the protocol’s functionality, their computational cost is significantly lower than modular exponentiations and is thus neglected in this dominant cost analysis.
At the Data Collector: DC’s primary computational burden occurs in the final decryption and extraction step. DC receives the final aggregated ciphertext, $\sigma^{\prime}_{\mathrm{sum},k}$, which consists of two Paillier ciphertexts $(C_{\mathrm{final},R},C_{\mathrm{final},I})$. DC performs 2 Paillier decryptions (one for the real part and one for the imaginary part). Each Paillier decryption primarily involves one modular exponentiation modulo $n^{2}$. Therefore, the DC performs approximately 2 modular exponentiations modulo $n^{2}$ per round.
Network-Wide Aggregation (Aggregator Nodes): The aggregation tree involves N1N-1 aggregation steps in total (for $N$ sensors in a binary tree). Each aggregation step involves two modular multiplications modulo $n^{2}$ (one for the real component sum and one for the imaginary component sum). These operations, while numerous, are computationally less intensive than modular exponentiations and are therefore not included in this dominant cost count.
The computational overhead for each party during a single data exchange round is analyzed based on the number of dominant modular exponentiations. The Sensor ($\mathrm{S}_{j}$) performs 4 integer modular exponentiations modulo $n^{2}$ for encrypting its watermarked data (two for the real part and two for the imaginary part). In contrast, the Data Collector (DC) incurs a comparatively lighter computational burden, totaling approximately 2 integer modular exponentiations modulo $n^{2}$ per round for decrypting the aggregated ciphertext. The network-wide aggregation process, while involving numerous modular multiplications, is not included in this dominant cost count as these operations are less intensive than exponentiations. This distribution of workload, where the Sensor bears a higher computational burden in terms of dominant Paillier exponentiations, is an important consideration for resource-constrained sensor deployments, implying the need for adequate computational capabilities at the sensor level.
On the other hand, the communication complexity is measured by the number of messages exchanged in a single data exchange round ($t_{k}$).
  • DC to Sensors ($\lambda_{k}$ distribution): The DC sends the symmetrically encrypted challenge factor $\mathrm{Enc}_{K_{\mathrm{sym},j}}(\lambda_{k})$ to each of the $N$ sensors. This accounts for $N$ messages from DC to sensors.
  • Sensor-to-Sensor and Sensor-to-DC (Aggregation): The cascading summation process in a binary tree structure involves messages being passed up the tree. For instance, if $N$ is a power of 2, the first level of aggregation generates $N/2$ messages, the second level $N/4$, and so on, until the last level generates 1 message. This sums up to $N/2+N/4+\ldots+1=N-1$ messages for the intermediate aggregation steps. A final message is then sent from the root aggregator sensor ($\mathrm{S}_{N}$) to the DC. Thus, the total number of messages for the aggregation phase (from sensors to DC) is (N1)+1=N(N-1)+1=N messages, regardless of whether $N$ is a power of 2.
Combining the DC-to-sensor and the aggregation messages, the overall number of messages exchanged in the network per round is $N+N=2N$ messages. This indicates an efficient communication pattern, as the total message count scales linearly with the number of sensors.
6 Formal Security and Privacy Analysis
In this section, we present to a formal theorem-proof evaluation of the proposed protocols. We demonstrate the resilience of the H[i][i]dden framework against the attack model defined in Section 5.1.3 through rigorous cryptographic proofs.
6.1 Confidentiality and Privacy Proofs
Theorem 1 (Resistance to Eavesdropping).
The H[i][i]dden framework ensures the confidentiality of $d_{k}$ (or $d_{j,k}$), $w_{k}$, and $\lambda_{k}$ against a passive eavesdropper $\mathcal{A}$ under the assumption that the Gaussian ElGamal and Paillier cryptosystems are IND-CPA (Indistinguishability under Chosen Plaintext Attack) secure.
Proof.
Suppose there exists an adversary $\mathcal{A}$ that can recover any part of $d_{k}$ (or $d_{j,k}$), $w_{k}$, or $\lambda_{k}$ from the transmitted ciphertexts with non-negligible advantage. We can construct a simulator $\mathcal{B}$ that uses $\mathcal{A}$ to break the semantic security (IND-CPA) of the underlying cryptosystems.
[…]
In H[i][i]dden-AggP, the individual watermarked data points $\delta_{j,k}^{\prime}$ are protected by the IND-CPA security of the Paillier cryptosystem. The challenge factor $\lambda_{k}$ is transmitted to each sensor $\mathrm{S}_{j}$ using symmetric encryption with a unique key $K_{\mathrm{sym},i}$, whose secrecy is established during initialization via Paillier encryption. If $\mathcal{A}$ could recover $\lambda_{k}$ or any component of the sensor data, the simulator $\mathcal{B}$ would be able to break the underlying semantic security of either the Paillier scheme or the symmetric cipher. Given the assumed security of these primitives, no such adversary $\mathcal{A}$ can exist, and confidentiality is maintained. ∎
Theorem 2 (Aggregation Privacy).
The H[i][i]dden-AggP protocol preserves the privacy of individual sensor readings $d_{j,k}$ against honest-but-curious partial aggregators $\mathrm{S}_{\ell}$.
Proof.
In H[i][i]dden-AggP, a partial aggregator $\mathrm{S}_{\ell}$ receives pairs of ciphertexts $(c_{j,k,R},c_{j,k,I})$, where $c_{j,k,R}=\mathrm{Enc}_{\mathsf{K_{pub}}}(\operatorname{Re}(\delta^{\prime}_{j,k}),r_{j,k,R})$ and $c_{j,k,I}=\mathrm{Enc}_{\mathsf{K_{pub}}}(\operatorname{Im}(\delta^{\prime}_{j,k}),r_{j,k,I})$. The aggregator computes the product component-wise: $C_{\mathrm{sum},R}=\prod_{j}c_{j,k,R}\,(\operatorname{\mathrm{mod}}n^{2})$ and $C_{\mathrm{sum},I}=\prod_{j}c_{j,k,I}\,(\operatorname{\mathrm{mod}}n^{2})$. Due to the IND-CPA security of the Paillier cryptosystem, the aggregator cannot distinguish between the encryption of a valid component (real or imaginary) and an encryption of a random value in $\mathbb{Z}_{n}$. Since $\mathrm{S}_{\ell}$ does not possess the private key $\mathsf{K_{priv}}$, the view of $\mathrm{S}_{\ell}$ consists solely of pairs of ciphertexts which reveal zero information about the underlying component values. Consequently, the aggregator is unable to learn any information about the Gaussian integer $\delta^{\prime}_{j,k}=\lambda_{k}(d_{j,k}+iw_{k})$, ensuring that individual readings $d_{j,k}$ remain private during the aggregation process. ∎
6.2 Integrity, Provenance, and Active Attack Proofs
Theorem 3 (Resistance to Replay Attacks).
An active adversary $\mathcal{A}$ cannot successfully replay a ciphertext (or a component thereof) captured at time $t_{k^{\prime}}$ to be accepted at time $t_{k}$ (where $k\neq k^{\prime}$) with non-negligible probability.
Proof.
At each time sample $t_{k}$, the Data Collector expects a specific Gaussian integer result $\delta^{\prime}_{k}$ (or $\sigma^{\prime}_{k}$) that, upon extraction with the current secret challenge factor $\lambda_{k}$, yields the expected time-dependent watermark $w_{k}$ (or $Nw_{k}$ in the aggregated case). We distinguish two cases:
  1. 1.
    Individual case (H[i][i]dden-EG): If $\mathcal{A}$ replays a ciphertext from $t_{k^{\prime}}$ at time $t_{k}$, the DC recovers $\delta^{\prime}_{k^{\prime}}=\lambda_{k^{\prime}}\delta_{k^{\prime}}$. The extraction yields:
    $\tilde{\delta}_{k}=\frac{\delta^{\prime}_{k^{\prime}}}{\lambda_{k}}=\frac{\lambda_{k^{\prime}}}{\lambda_{k}}(d_{k^{\prime}}+iw_{k^{\prime}}).$
    Since $\lambda_{k}$ and $\lambda_{k^{\prime}}$ are secret and refreshed every round, the resulting ratio $\lambda_{k^{\prime}}/\lambda_{k}$ is unknown to $\mathcal{A}$, making $\operatorname{Im}(\tilde{\delta}_{k})$ effectively random. The condition $\operatorname{Im}(\tilde{\delta}_{k})=w_{k}$ is satisfied only with probability $2^{-B}$.
  2. 2.
    Aggretated case (H[i][i]dden-AggP): If $\mathcal{A}$ captures a sensor’s component ciphertexts $(c_{j,k,R},c_{j,k,I})$ at $t_{k^{\prime}}$ and replays them into the aggregate for $t_{k}$, the decrypted aggregate sum is:
    $\tilde{\sigma}_{k}=\left(\sum_{l\neq j}\lambda_{k}\delta_{l,k}\right)+\lambda_{k^{\prime}}\delta_{j,k^{\prime}}.$
    The DC extracts the aggregate watermark:
    $\operatorname{Im}(\tilde{\sigma}_{k}/\lambda_{k})=(N-1)w_{k}+\operatorname{Im}\left(\frac{\lambda_{k^{\prime}}}{\lambda_{k}}\delta_{j,k^{\prime}}\right).$
    For the aggregate to be accepted, the condition $\operatorname{Im}(\tilde{\sigma_{k}}/\lambda_{k})=Nw_{k}$ must hold, implying that
    $\operatorname{Im}\left(\frac{\lambda_{k^{\prime}}}{\lambda_{k}}(d_{j,k^{\prime}}+iw_{k^{\prime}})\right)=w_{k}.$
    The probability of an accidental collision where the resulting imaginary part matches the current expected watermark $w_{k}$ is exactly $2^{-B}$.
In both cases, the attacker must satisfy a condition involving three independent, secret, and time-varying parameters: the old watermark $w_{k^{\prime}}$, the old challenge $\lambda_{k^{\prime}}$, and the new challenge $\lambda_{k}$. Thus, detection occurs with probability $1-2^{-B}$ in either scenario. ∎
Theorem 4 (Resistance to Man-in-the-Middle (MitM) Attacks).
Any unauthorized modification of a legitimate ciphertext by $\mathcal{A}$ will be detected by the DC with overwhelming probability, preserving data integrity.
Proof.
In the H[i][i]dden framework, the DC chooses a modulus $p$ (for H[i][i]dden-EG) or $n$ (for H[i][i]dden-AggP) sufficiently large to ensure exact recovery of the Gaussian integer value $\delta^{\prime}_{k}=\lambda_{k}\delta_{k}$ or the aggregate $\sigma^{\prime}_{k}=\lambda_{k}\sigma_{k}$, respectively, upon decryption of a valid ciphertext. Consider a MitM adversary $\mathcal{A}$ who intercepts a ciphertext of $\delta^{\prime}_{k}$ (for H[i][i]dden-EG) or $\delta^{\prime}_{j,k}$ (for H[i][i]dden-AggP) and transmits a modified version, resulting in an induced error $\epsilon=e_{R}+ie_{I}$ in the plaintext domain after decryption.
  1. 1.
    Individual case (H[i][i]dden-EG): The DC extracts:
    $\tilde{\delta}_{k}=\frac{\delta^{\prime}_{k}+\epsilon}{\lambda_{k}}=\delta_{k}+\frac{\epsilon}{\lambda_{k}}.$
    Expanding using $\lambda_{k}^{-1}=(a_{k}-ib_{k})/(a_{k}^{2}+b_{k}^{2})$:
    $\tilde{\delta}_{k}=(d_{k}+iw_{k})+\frac{(e_{R}a_{k}+e_{I}b_{k})+i(e_{I}a_{k}-e_{R}b_{k})}{a_{k}^{2}+b_{k}^{2}}.$
    The extracted watermark is (imaginary part):
    $\tilde{w}_{k}=w_{k}+\frac{e_{I}a_{k}-e_{R}b_{k}}{a_{k}^{2}+b_{k}^{2}}.$
  2. 2.
    Aggregated case (H[i][i]dden-AggP): The DC extracts the aggregate
    $\tilde{\sigma}_{k}=\frac{\sigma^{\prime}_{k}+\epsilon}{\lambda_{k}}=\sigma_{k}+\frac{\epsilon}{\lambda_{k}}.$
    Since $\sigma_{k}=S_{k}+iNw_{k}$ (Section 3.3):
    $\tilde{\sigma}_{k}=(S_{k}+iNw_{k})+\frac{(e_{R}a_{k}+e_{I}b_{k})+i(e_{I}a_{k}-e_{R}b_{k})}{a_{k}^{2}+b_{k}^{2}}.$
    The extracted watermark is (imaginary part divided by $N$):
    $\tilde{w}_{k}=w_{k}+\frac{e_{I}a_{k}-e_{R}b_{k}}{N(a_{k}^{2}+b_{k}^{2})}.$
In both cases, for the modification to go undetected, the induced imaginary shift must be zero, implying $e_{I}a_{k}-e_{R}b_{k}=0$. Since $\mathcal{A}$ blindly manipulates ciphertexts without knowledge of $(a_{k},b_{k})$, the attacker cannot choose a modification that purposely aligns $\epsilon$ with the secret “slope” of $\lambda_{k}$. For a large secret space, the probability of an accidental collision is negligible, ensuring protection against MitM modifications. ∎
Theorem 5 (Resistance to Masquerading and False Data Injection (FDI)).
The H[i][i]dden framework ensures resistance to masquerading and FDI attacks; a malicious entity $\mathcal{A}$ cannot successfully inject fraudulent readings $d_{\mathrm{fake}}$ that the DC accepts as originating from a legitimate sensor $\mathrm{S}_{j}$.
Proof.
An FDI attack differs from MitM modification in that $\mathcal{A}$ attempts to inject an entirely synthetic reading $d_{\mathrm{fake}}$ without access to a valid message’s ciphertext. To succeed, $\mathcal{A}$ must generate a ciphertext that decrypts to a specific target value $\delta^{\prime}_{\mathrm{target}}=\lambda_{k}(d_{\mathrm{fake}}+iw_{k})$.
As established in the Attack Model, $\mathcal{A}$ does not possess the secret challenge factor $\lambda_{k}$ nor the current watermark $w_{k}$. In the H[i][i]dden-EG protocol, $\lambda_{k}$ is transmitted to the sensor $\mathrm{S}_{j}$ encrypted using the DC’s public key; its confidentiality is thus guaranteed by the semantic security of the Gaussian ElGamal scheme. Similarly, in the H[i][i]dden-AggP protocol, $\lambda_{k}$ is protected by a unique symmetric key $K_{\mathrm{sym},j}$ shared between the DC and $\mathrm{S}_{j}$.
Consequently, $\mathcal{A}$ is unable to learn $\lambda_{k}$ by eavesdropping in either protocol. Without these secrets, $\mathcal{A}$ would need to guess both the specific multiplicative factor $\lambda_{k}$ and the expected watermark $w_{k}$ for the current time frame. If the secret space for $(a_{k},b_{k})$ has cardinality $|S_{\lambda}|$ and the watermark space has cardinality $2^{B}$, the probability of $\mathcal{A}$ successfully crafting an injected value that survives the DC’s division and watermark extraction test is approximately:
$P(\mathrm{success})\approx\frac{1}{|S_{\lambda}|}\cdot\frac{1}{2^{B}}.$
For standard cryptographic parameters, this probability is negligible. Thus, the successful extraction of $w_{k}$ serves as a proof of origin and authenticity, fulfilling the requirement for data provenance. ∎
Theorem 6 (Perfect Reversibility and Accuracy).
The H[i][i]dden framework provides zero-distortion reconstruction of the host data if the range bounds $\max(|\operatorname{Re}(\delta^{\prime}_{k})|,|\operatorname{Im}(\delta^{\prime}_{k})|)<p/2$ (for H[i][i]dden-EG) and $\max(|\operatorname{Re}(\sigma^{\prime}_{k})|,|\operatorname{Im}(\sigma^{\prime}_{k})|)<n/2$ (for H[i][i]dden-AggP) are satisfied.
Proof.
The correctness of the reconstruction relies on the modular decryption recovering the exact Gaussian integer result of the watermarking process:
  1. 1.
    Individual case (H[i][i]dden-EG): Decryption modulo $p$ in a centered representative system yields the exact value $\delta^{\prime}_{k}=\lambda_{k}(d_{k}+iw_{k})$ if its real and imaginary parts satisfy $|\operatorname{Re}(\delta^{\prime}_{k})|<p/2$ and $|\operatorname{Im}(\delta^{\prime}_{k})|<p/2$.
  2. 2.
    Aggregated case (H[i][i]dden-AggP): Component-wise decryption modulo $n$ recovers the aggregate components exactly if $|\operatorname{Re}(\sigma^{\prime}_{k})|<n/2$ and $|\operatorname{Im}(\sigma^{\prime}_{k})|<n/2$.
In both scenarios, once the exact Gaussian integer ($\delta^{\prime}_{k}$ or $\sigma^{\prime}_{k}$) is recovered, standard complex division by the current secret challenge factor $\lambda_{k}$ in the field $\mathbb{C}$ yields the exact watermarked data $\delta_{k}=d_{k}+iw_{k}$ or the aggregate $\sigma_{k}=S_{k}+iNw_{k}$. The host data ($d_{k}$ or $S_{k}$) is then perfectly reconstructed by taking the real part of the result, ensuring zero distortion. ∎
In short, the formal evaluation conducted through Theorems 16 provides a comprehensive verification of the system requirements and attack model established in Section 5.1. The robustness against eavesdropping (Theorem 1) and the guarantee of aggregation privacy (Theorem 2) satisfy the requirements of confidentiality and privacy. The probabilistic detection of replay (Theorem 3) and the algebraic detection of MitM modifications (Theorem 4) fulfill the integrity requirement. Furthermore, the analysis of masquerading and FDI resistance (Theorem 5) validates the data provenance and authenticity properties by binding successful watermark extraction to the shared secrets. Finally, the accuracy proof (Theorem 6) confirms that the protocols achieve perfect reversibility under defined conditions. Consequently, the H[i][i]dden framework is formally proven to be secure against both passive and active adversarial capabilities $\mathcal{A}$ within the distributed environment.
7 Conclusion
This paper introduces Hiding in the Imaginary Domain with Data Encryption (H[i][i]dden), a novel framework for RDH that operates uniquely within the complex domain. By intrinsically mixing data and watermark through complex number arithmetic, the H[i][i]dden approach offers perfect reversibility and allows for embedding a watermark of in principle unlimited size, directly addressing key limitations of existing RDH techniques for individual and dispersed data. Two distinct protocols embodying this framework are proposed: the H[i][i]dden-EG protocol, providing joint RDH and encryption for individual data points by extending ElGamal encryption to Gaussian integers, and the H[i][i]dden-AggP protocol, which enables privacy-preserving data aggregation by taking advantage of a novel component-wise application of Paillier encryption to complex-valued data.
The proposed H[i][i]dden framework and its instantiations offer novel solutions for ensuring data integrity and authenticity in scenarios involving sensitive, distributed, and potentially resource-constrained environments like Wireless Sensor Networks and IoT ecosystems. By integrating advanced cryptographic primitives with innovative complex-domain embedding, the proposed methods provide strong security guarantees beyond simple data hiding, including confidentiality and privacy-preserving aggregation. This work demonstrates a significant step towards building more trustworthy data pipelines from diverse sources, where the provenance and integrity of information are strong requirements.
As directions for future work, one key area involves exploring the inclusion of other homomorphic encryption schemes, such as BGN or FHE schemes, to operate efficiently within the complex domain or over Gaussian integers. Such extensions could unlock new homomorphic capabilities (e.g., multiple multiplications on ciphertexts) or improved efficiency for complex-valued computations. In addition, this could provide post-quantum security, an aspect that is not covered by ElGamal or Paillier.
Another future research direction is the potential application of the H[i][i]dden approach to other domains, such as media information or non-dispersed time series data. By adapting integer-based media formats to the complex domain for embedding, this method could offer a novel approach to traditional image, audio, or video watermarking challenges, taking advantage of the inherent mixing properties and high capacity of the H[i][i]dden scheme.
Acknowledgments
This work has been supported, in part, by the “SAFE” (PID2024-156914OB-C41) and “SECURING” (PID2021-125962OB-C31) grants, funded by the Ministry of Science and Innovation thorugh the AEI and the European Regional Development Fund (ERDF), and by the ARTEMISA International Chair of Cybersecurity (C057/23) and the DANGER Strategic Project of Cybersecurity (C062/23), both funded by the Spanish National Institute of Cybersecurity through the European Union—NextGenerationEU and the Recovery, Transformation, and Resilience Plan.
Appendix A Modular Arithmetic over Gaussian Integers
It is worth pointing out that Gaussian prime numbers have some differences compared to the regular integer case. The obvious one is that Gaussian primes can have both real and imaginary parts. The second main difference is that not all integer primes are also Gaussian primes. For example, prime factorizations exists for both 2 and 5 in $\mathbb{Z}[i]$: 2=(1+i)(1i)2=(1+i)(1-i) and 5=(2+i)(2i)5=(2+i)(2-i). The integer primes $p$ that are also Gaussian primes satisfy the following condition $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$ as proven, for example, in (Conrad, 2016).
Figure 4: Representatives of Gaussian integers $(\operatorname{\mathrm{mod}}7)$, or $\mathbb{Z}[i]_{7}$.
Regarding modular arithmetic, we need to extend the definition of integer division to the complex domain. For general Gaussian integers $\alpha,\beta\in\mathbb{Z}[i]$, the Gaussian integer division is defined as $\alpha=\kappa\beta+\rho$, for some residual $\rho$ with $\mathrm{N}(\rho)<\mathrm{N}(\beta)$, where $\mathrm{N}(\cdot)$ stands for the norm of a complex number, which is defined as follows:
$\mathrm{N}(a+ib)=\left|a+ib\right|^{2}=a^{2}+b^{2}.$
This definition preserves the intuition of integer division for Gaussian integers since, for each number $\alpha$ and each divider $\beta$ in Gaussian integers, a quotient $\kappa$ and a remainder $\rho$ can be found such that $\alpha=\kappa\beta+\rho$, and $\rho$ is, in some sense, smaller than the divider $\beta$. Here, the norm of complex number is the criterion for determining whether a complex number is smaller than another.
With this definition of the integer division, we can apply the same rules of modular arithmetic as for standard integers, i.e., for $\alpha,\beta,\kappa,\rho\in\mathbb{Z}[i]$, if $\alpha=\kappa\beta+\rho$, we have $\alpha\equiv\rho\,(\operatorname{\mathrm{mod}}\beta)$, since $\beta|(\alpha-\rho)$.
It is well known that, in regular integers, for a prime number $p\in\mathbb{Z}$, the elements of the cyclic group $\mathbb{Z}_{p}$ (or $\mathbb{Z}/p\mathbb{Z}$) are $0,1,\dots,p-1$, while the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$ has the same elements, excluding 0. The order of the corresponding cyclic groups $\mathbb{Z}[i]_{\pi}$ and ${\mathbb{Z}[i]}^{*}_{\pi}$ is not so straightforward when $\pi$ is a Gaussian prime. For Gaussian integers, given a prime number $\pi\in\mathbb{Z}[i]$, the norm $\mathrm{N}(\pi)$ determines the order of the additive cyclic group $\mathbb{Z}[i]_{\pi}$, whereas $\mathrm{N}(\pi)-1$ is the order of the multiplicative cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ (because of the exclusion of 0). Proofs for these facts can also be found in (Conrad, 2016).
As a practical example, the selection of representatives for the cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ for $\pi=7$, which is both an integer and a Gaussian prime, is discussed below. First, we need to take into account that many possible selections of representatives exists, similar to the integer case. For example, 13 and 6 refer to the same equivalence class in both ${\mathbb{Z}}^{*}_{7}$ and ${\mathbb{Z}[i]}^{*}_{7}$, since 13 and 6 differ in a multiple of 7 (in fact, exactly 7). The only relevant consideration for the selection is not to include more than one representative for the same class. A method for such a selection is described in (Conrad, 2016) both when the divider has real and imaginary parts different from zero, and when the imaginary part is zero (the case of the Gaussian prime $\pi=7$).
The procedure is illustrated in Fig. 4: We take $7$ (the divider) and $7i$ (by multiplying by $i$) and draw a square which has 7 and $7i$ as the left and lower edges (shown in green color in the picture). The square is completed with the purple edges that connect 7 and $7i$ with the vertex $7+7i$. The representatives of $\mathbb{Z}[i]_{7}$ are the Gaussian integers that lie within the square and the ones which are on the green edges, but excluding the ones on the purple edges, since they represent the same equivalence class as the elements on the green ones.
Figure 5: “Canonical” epresentatives of Gaussian integers $(\operatorname{\mathrm{mod}}7)$, or $\mathbb{Z}[i]_{7}$.
With this particular selection, the set of values of the cyclic multiplicative group ${\mathbb{Z}[i]}^{*}_{7}$ can be easily represented as
${\mathbb{Z}[i]}^{*}_{7}=\left\{x+iy\mid x,y\in\{0,1,\dots,6\right\}\text{ and }(x,y)\neq(0,0)\},$
with no equivalent elements. This leads to the following observations:
  • The order of the additive cyclic group ${\mathbb{Z}[i]}_{p}$, for $p\in\mathbb{Z}$ (i.e., $p$ is a regular integer) is $p^{2}$.
  • If $p\in\mathbb{Z}$ is both a regular prime and a Gaussian prime, i.e., when $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$, then ${\mathbb{Z}[i]}^{*}_{p}$ is a multiplicative cyclic group of order $p^{2}-1$.
This set of representatives is particularly convenient to carry out modular arithmetic operations, since $(a+ib)\,(\operatorname{\mathrm{mod}}7)$ is simply $a\,(\operatorname{\mathrm{mod}}7)+ib\,(\operatorname{\mathrm{mod}}7)$. This means that we can just apply the modular operator to the real and the imaginary parts separately, which is easy to handle with most of the existing programming languages.
This selection of representatives is not “canonical”, in the sense that it includes some elements whose norms are greater than $\mathrm{N}(7)=7^{2}=49$. For example $\mathrm{N}(5+5i)=5^{2}+5^{2}=50$. A more “canonical” set of representatives for ${\mathbb{Z}[i]}^{*}_{7}$ would be the following (Fig. 5):
${\mathbb{Z}[i]}^{*}_{7}=\left\{x+iy\mid x,y\in\{-3,-2,\dots,3\}\text{ and }(x,y)\neq(0,0)\right\},$
which is obtained by shifting the values of Fig. 4 by 3-3 units in each axis. This new selection of representatives is intuitively closer to that usually made for the cyclic groups defined using standard integers, but the one shown in Fig. 4 is more convenient for computational operations. In any case, both are completely equivalent. For example, $5+5i\equiv-2-2i\,(\operatorname{\mathrm{mod}}7)$, where the former is one of the points represented in Fig. 4 and latter only appears in Fig. 5. Note that the norm of 22i-2-2i is $2^{2}+2^{2}=8$, which, of course, is lower than $N(7)=49$. Hence, although $5+5i$ has a norm greater than $\mathrm{N}(7)$, an equivalent value can easily be found with a norm below that of the modulus. The norms of all values represented in Fig. 5 are lower than 49.
Remark 18.
When the data to be represented is signed (including samples with both positive and negative components), the representation of Fig. 5 is mandatory. In such a case, a simple shift can map every value of Fig. 4 to its corresponding value of Fig. 5 (or conversely). $\triangle$